### v1

# A Comprehensive Evaluation on Quantization Techniques for Large Language Models

Abstract
For large language models (LLMs), post-training quantization (PTQ) can significantly reduce memory footprint and computational overhead. Model quantization is a rapidly evolving research field. Though many papers have reported breakthrough performance, they may not conduct experiments on the same ground since one quantization method solution usually contains multiple components. In addition, it is also important to analyze the theoretical connections among numerous existing methods to gain in-depth understanding in this field. To bridge these gaps, in this paper, we timely conduct an extensive review of the state-of-the-art methods and also perform comprehensive evaluations on the same ground to ensure fair comparisons of these methods. To our knowledge, this fair and extensive investigation remains critically important yet significantly underexplored in the research community.
To better understand the theoretical connections, we decouple the published quantization methods into two steps: pre-quantization transformation and quantization error mitigation. We define the former as a preprocessing step applied before quantization to reduce the impact of outliers, making the data distribution flatter and more suitable for quantization. Quantization error mitigation, on the other hand, involves techniques that offset the errors introduced during quantization, thereby enhancing model performance. Meanwhile, we extensively evaluate and analyze the impact of different components of quantization methods. Additionally, we also timely analyze and evaluate the latest MXFP4 data format and its performance. Our experimental results demonstrate that the optimized rotation and scaling yield the best performance for pre-quantization transformation, and combining low-rank compensation with GPTQ occasionally outperforms using GPTQ alone for quantization error mitigation. Furthermore, we explore the potential of the latest MXFP4 quantization and reveal that the optimal pre-quantization transformation strategy for INT4 does not generalize well to MXFP4, inspiring further investigation in this field.
keywords:
Model Quantization, Post-training Quantization , LLM
1 Introduction
Large language models (LLMs) have achieved remarkable performance on diverse tasks and are increasingly being applied in various scenarios. Nonetheless, LLMs inference process incurs significant memory and computational overhead, leading to high inference latency and posing challenges for computations on source-limited GPUs.
Post-training quantization (PTQ) is a promising solution to reduce the memory and computational cost of LLMs, enhancing their inference efficiency. LLMs quantization has been studied in a wide spectrum of precision configurations. Representative settings include W4A4 [1, 2, 3, 4], W8A8 [5, 6, 7] and W4A16 [8, 9, 10, 11]: W4A4 quantizes both weights and activations to 4-bit precision (INT4); W8A8 and W4A16 quantize weights to INT8 and INT4 and activations to INT8 and FP16/BF16 respectively. Among these 3 settings, W4A4 is the current research focus because it provides the best performance in terms of both memory efficiency and inference speed, with the greatest research significance.
However, despite the wide range of methods explored in previous studies, W4A4 quantization still struggles to achieve lossless performance, especially for smaller LLMs, such as those with 1B, 3B, and 7B parameters, which often results in increased perplexity and degraded performance on various tasks. Furthermore, LLaMA3 experiences a larger performance drop after quantization compared to LLaMA2 when using many popular quantization methods [3, 4, 12], facing a greater demand for better W4A4 quantization methods. Therefore, W4A4 quantization for LLMs still remains a challenging task.
To review and evaluate the state-of-the-art quantization methods, in particular, under the setting of W4A4, we find two main characteristics of these methods. First, existing methods are often combinations of multiple techniques, since using a single technique alone typically fails to achieve excellent quantization performance. In SpinQuant [3], after applying trained rotations to adjust the model, the GPTQ [8], a gradient-based quantization error mitigation method is further applied to quantize the rotated weights. Similarly, OSTQuant [8] employs multiple techniques to adjust the weights and activations, including rotation and scaling. At the same time, there are many weight-only quantization methods, some of which can be combined with pre-quantization transformation techniques and extended to weight-activation quantization. AWQ [11] shares similar pre-quantization transformation (scaling) with SmoothQuant [6], even though their motivations differ. Also, QuIP# [10] shares similar pre-quantization transformation (random hadamard rotation) with QuaRot [13]. Apart from the methods that involve both weight and activation transformations, there are also approaches that are orthogonal to pre-quantization transformation that can be used to compensate for the loss between quantized models and full precision models, such as GPTQ [8] and low-rank branches computed from quantization error [14, 15, 16]. Second, quantization involves a variety of configurations, such as granularity and symmetric or not. Different settings can lead to different performance outcomes and often involve trade-offs between efficiency and accuracy. For different precision levels, the optimal quantization settings may vary. Evaluating the performance under various quantization configurations for W4A4 quantization is of great importance.
In this paper, as aforementioned in the first characteristic, the existing quantization solutions are very complicated and they usually consists of many algorithms. To make our evaluation on the same ground and achieve a better understanding of existing methods, we decompose the quantization process into two key steps: pre-quantization transformation and quantization error mitigation. The pre-quantization transformation step refers to preprocessing the data before quantization in order to mitigate the effect of outliers making the data distribution smoother and more quantization-friendly. The quantization error mitigation step involves applying techniques to compensate for the errors introduced by quantization, thereby reducing the overall quantization error and improving performance. For each step, we select several state-of-the-art techniques and conduct an evaluation of the effectiveness of their combinations. Second, existing solutions usually do not use exactly the same setting, e.g. granularity, leading to unfair comparison. To bridge this gap, we conduct a comprehensive evaluation on the same quantization configurations and provide recommendations on selecting appropriate configurations. And we integrate the methods used in this paper in our code library to facilitate performance evaluation.
Furthermore, the NVIDIA’s GeForce RTX 50 Series latestly released based on the new RTX Blackwell architecture includes support for the FP4 data format, including NVFP4 and MXFP4, in its new Tensor Cores, which show great advantages over INT4. However, the research on FP4 has not been well explored. To bridge this gap, we conduct research on weight-activation FP4 quantization using the popular methods in INT4, showing the great timeliness of our research.
Methods Data Normalization Error Compensation
SmoothQuant [6] scale -
Outlier Suppression+ [7] scaling + shifting -
GPTQ [8] - GPTQ
ZeroQuant-v2 [14] - low-rank
QuIP [9] rotation GPTQ
AWQ [11] scaling -
Atom [1] reorder GPTQ
OmniQuant [2] optimized scaling + shifting GPTQ
Qserve [17] scaling + rotation GPTQ
LQER [18] - low-rank
QuaRot [13] rotation GPTQ
SpinQuant [3] optimized rotation GPTQ
ResQ [19] reorder + rotation GPTQ
FlatQuant [12] optimized scaling + rotation GPTQ
OSTQuant [4] optimized scaling + rotation GPTQ
Table 1: Two steps decomposition of quantization methods.
2 Review on Existing Quantization Methods
In this section, we first review recent quantization algorithms in Section 2.1 and Section 2.2 in terms of two steps: pre-quantization transformation and quantization error mitigation. Table 1 shows our decomposition of the popular quantization methods. Section 2.3 and Section 2.4 focus on settings, including symmetricity and granularity of quantization methods. Meanwhile, Table 2 summarizes the quantization settings reported in the experimental results of these methods.
Methods Precision Symmetric Granularity
W A
SmoothQuant [6] W8A8 T F T,C
GPTQ [8] W3A16,W4A16 T - G
ZeroQuant-v2 [14] W4A16,W3A16,W2A16,W4A8 T,F T,F G
QuIP [9] W2A16,W3A16,W4A16 T - G
AWQ [11] W3A16,W4A16 T - G
Atom [1] W3A3,W4A4 T F G
OmniQuant [2] W2A16,W3A16,W4A16,W4A4,W6A6,W8A8 F F C
Qserve [17] W4A8 T F C
LQER [18] W4A6,W4A8 T F G
QuaRot [13] W4A4 T F C,G
SpinQuant [3] W4A4,W4A8 T F C
ResQ [19] W4A4 T F C
FlatQuant [12] W4A4 T F C
OSTQuant [4] W4A4 T F C
Table 2: Symmetry and granularity of quantization methods. In granularity, T stands for per-tensor, C represents per-channel or per-token, and G means per-group
2.1 Pre-quantization transformation
Pre-quantization transformation is a step before quantization, applying operations such as shifting, scaling, and rotation to process weights and activations. These transformations aim to flatten and smooth the hard-to-quantize tensors, thereby enhancing the quantization performance. LLMs are difficult to quantize due to the outliers in the activations. Compared to most of the activation values, the magnitude of outliers is substantially higher, often exceeding 100x larger in some layers. The outliers will dominate the maximum magnitude measurement, leading to low effective quantization levels. This pre-quantization transformation step is illustrated in Fig. 1.
Figure 1: Pre-quantization transformation: Transforming activations and weights to render them more easier to quant
2.1.1 Shifting
Activations distribution is often asymmetric, and this asymmetry varies across different channels. Aligning the center of each channel can improve the performance of quantization, since we quantize activations per-token rather than along the channel dimension. Outlier Suppression+ [7] proposes shifting activation values to eliminate the asymmetry feature across channels. The shifting process is to use a vector that transfer the channels of the activation. For a linear layer, it is formulated as:
$Y=XW+B=(\underbrace{(X-T)}_{\hat{X}})W+(\underbrace{TW+B}_{\hat{B}})$ (1)
where the layer input $X\in\mathbb{R}^{T\times C_{in}}$ is shifting and becomes $\hat{X}$. $T\in\mathbb{R}^{1\times C_{in}}$ is a vector representing the channel-wise shifting parameters. To keep mathematical equivalent, an additional term, the product of $T$ and $W\in\mathbb{R}^{C_{in}\times C_{out}}$, is added to the bias $B\in\mathbb{R}^{1\times C_{out}}$. Outlier Suppression+ [7] uses calibration to get the vector $T$, which is computed channel-wise as the average of its minimum and maximum values:
$t_{i}=\frac{\min(X_{i})+\max(X_{i})}{2}$ (2)
where $t_{j}$ denotes the $j$-th element of $T$, corresponding to the $j$-th channel. $X_{j}$ represents the $j$-th input channel of $X$. OmniQuant [2] employs learnable channel-wise shifting for outlier suppression, combined with the learnable scaling technique discussed in Section 2.1.2, and jointly optimize these two kinds of pre-quantization transformation parameters.
2.1.2 Scaling
Scaling is a commonly used pre-quantization transformation method, which scales activations in different input channels before quantization, and inversely scales the corresponding weight channels to make the results stay unchanged. SmoothQuant [6] first propose using scaling to smooth the quantization difficulty from activations to weights. Compared to weights, activations have more outliers that make quantization harder. It is observed that outliers in activations appear in certain channels, so applying channel-wise scaling to activations can eliminate the outliers to a certain degree, using the following equation for linear layers:
$Y=XW=(\underbrace{X\cdot S^{-1}}_{\hat{X}})(\underbrace{S\cdot W}_{\hat{W}})=\hat{X}\hat{W}$ (3)
where the weights of a linear layer $W$ and the layer input $X$ is scaling and becomes $\hat{X}$ and $\hat{W}$. $S$ is a diagonal matrix composed by scaling factors of different channels, which can be obtained in various ways. SmoothQuant [6] uses calibration to compute $S$. The calibration process can be formulated as:
$s_{j}=max(|X_{j}|)^{\alpha}/max(|W_{j}|)^{1-\alpha}$ (4)
where $X_{j}$ and $W_{j}$ denote the $j$-th input channel of $X$ and $W$, respectively. $s_{j}$ is the $j$-th diagonal element of $S$, corresponding to the $j$-th channel. $\alpha$ is a hyperparameter that controls the degree to which outliers are transferred between activations and weights.
AWQ [11], as a weight-only quantization method, approaches the problem from the perspective of protecting salient weights, applying activation-aware scaling. Unlike adopting the hand-crafting parameters in calibration, the scaling operation in OmniQuant [2] uses learnable scaling, where $S$ is learnable parameters that are optimized with block-wise quantization error minimization.
FlatQuant [12] and OSTQuant [4] both use learnable per-channel scaling, but at different stages of pre-quantization transformation, leading to different practical meanings. Both perform a diagonal scaling operation, yet FlatQuant scales before rotation, allowing its scaling factors to be initialized via Eq. (4), functionally similar to SmoothQuant. In contrast, OSTQuant applies scaling after rotation, meaning this diagonal scaling is a post-rotation adjustment rather than smoothing the original channels.
2.1.3 Rotation
The rotation operation is a technique that adjusts data distribution by multiplying the data matrix with an orthogonal matrix. This operation largely changes the activation distributions, reduces outliers in the data and thereby enhancing quantization accuracy. The rotation process in a linear layer can be formulated as:
$Y=XW=(\underbrace{XO}_{\hat{X}})(\underbrace{O^{\top}W}_{\hat{W}})=\hat{X}\hat{W}$ (5)
where the input $X$ is rotated with an orthogonal matrix $O$, resulting in $\hat{X}=XO$. To preserve the model’s output, the weight $W$ is simultaneously transformed by $O^{\top}$, which is the transpose of $O$.
QuIP [9] first introduced rotation with random orthogonal matrices to reduce incoherence in weight-only quantization. QuIP# [10] replaces this with random Hadamard matrices for greater efficiency. QuaRot [13] extends rotation to both weights and activations quantization using random Hadamard matrices, effectively eliminating outliers. SpinQuant [3] further improves stability by optimizing the rotation matrix on the Stiefel manifold to minimize quantization-induced performance variance, while still defaulting to random Hadamard for a few online rotations.
Rotation is a highly effective pre-quantization transformation technique for reducing outliers in low-bit quantization, and it has also been widely adopted in many other quantization methods (e.g., ResQ [19], QServe [17], FlatQuant [12] and OSTQuant [4], etc), which combine rotation with other pre-quantization transformation techniques to further improve performance.
2.2 Quantization error mitigation
The quantization error mitigation step employs techniques that mitigate the error introduced by quantization by compensating them. This error between the original model and the quantized model can be computed from calibration datasets, and be compensated to the original weights in linear layers or through low-rank branches. Fig. 2 demonstrates two commonly used quantization error mitigation methods.
Figure 2: Quantization error mitigation: compensate the error produced by quantization. (1) Self-compensation [8, 9] calculates weight compensation terms through quantization-induced loss errors; (2) Low-rank branch compensation [14, 16, 18, 15] constructs auxiliary compensation pathways via matrix factorization of post-quantization weight errors.
2.2.1 Self compensation
GPTQ [8] is a quantization method for weight quantization that uses Hessian-based quantization error mitigation. GPTQ builds on the OBQ [20] method that solves the error introduced by quantization and the optimal weight perturbation. For weight quantization, the loss error introduced by quantization can be defined as:
$\epsilon=\mathcal{L}(W_{q})-\mathcal{L}(W)=\mathcal{L}(W+\delta)-\mathcal{W}$ (6)
where $L(W)$ is the loss function of the model evaluated at weights $W$, and $\delta$ represents the weight perturbation. Using a Taylor expansion, we can get the Hessian-based approximation of the quantization loss, which can be formulated as:
$\epsilon=\nabla\mathcal{L}(W)^{\top}\delta+\frac{1}{2}\delta^{\top}H(W)\delta+\mathcal{O}(\left\|\delta\right\|^{3})\approx\frac{1}{2}\delta^{\top}H(W)\delta$ (7)
Unlike OBQ that uses greedy order to quantize weights, GPTQ quantized all rows of weights in a fixed order. The GPTQ algorithm performs quantization column by column in blocks of size B. For each column within the block, the quantization error is computed when quantized, and then compensates on the remaining weights. Once all columns of a block is quantized, the remaining weights in the matrix are also updated.
QuIP [9] also minimizes a quadratic proxy of the weight error to implement an optimal adaptive rounding procedure called LDLQ. In this method, an LDL decomposition of the Hessian matrix is performed to construct a feedback matrix $U$, which is then used to form a compensation term that corrects for the quantization-induced error. In QuIP, it is proven that for Eq. (7), performing an LDL decomposition on the Hessian matrix yields an optimal choice of $U$, and GPTQ is a special case of LDLQ.
2.2.2 Low-rank branch compensation
Beyond compensating errors directly on the original weights, an alternative approach is to introduce an auxiliary branch dedicated to error correction. Some prior works leverage a low-rank term to reconstruct quantization error (e.g., ZeroQuant-v2 [14], CALDERA [16], LQER [18], QERA [15], etc). After incorporating a low-rank branch, the output $Y$ is computed as:
$Y=XW=XW_{q}+X\Delta W\approx XW_{q}+XAB$ (8)
where $\Delta W=W-W_{q}$ denotes the weight reconstruction error introduced by quantization. $A\in\mathbb{R}^{C_{in}\times r}$, $B\in\mathbb{R}^{r\times C_{out}}$ are learnable low-rank matrices, and $r\ll\min(d,k)$ is the bottleneck rank. The product $AB\in\mathbb{R}^{d\times k}$ serves as a low-rank approximation of the quantization error.
ZeroQuant-v2 [14] and LQER [18] reconstruct the weight quantization error through SVD-based low rank approximation. LQER computes the low rank matrices $A,B$ by using SVD composition to the quantization error matrix $\Delta W$:
$\Delta W=U\Sigma V^{T}\approx\underbrace{U_{k}}_{A_{k}}\cdot\underbrace{\Sigma_{k}V_{k}^{T}}_{B_{k}}=A_{k}B_{k}$ (9)
where $U\in\mathbb{R}^{m\times n}$, $V\in\mathbb{R}^{n\times n}$, $\Sigma\in\mathbb{R}^{m\times n}$ and $U_{k}\in\mathbb{R}^{m\times k}$, $V_{k}\in\mathbb{R}^{n\times k}$, $\Sigma_{k}\in\mathbb{R}^{k\times k}$.
Based on LQER, L2QER, which is proposed in LQER, scales the $\Delta W$ before applying SVD, and inversely apply scales in $A_{k}$. This operation aims to more precisely approximated the quantization error corresponding to the salient weights. The process can be formulated as:
$\Delta\hat{W}=S\Delta W=U\Sigma V^{T}\approx\underbrace{U_{k}}_{A_{k}}\cdot\underbrace{\Sigma_{k}V_{k}^{T}}_{B_{k}}=A_{k}B_{k}$ (10)
$\hat{A_{k}}=S^{-1}A_{k},\hat{B_{k}}=Bk$ (11)
In addition to the aforementioned approaches, CALDERA [16] and QERA [15] provide analyses from different perspectives, deriving the optimal low-rank terms for quantization error reconstruction. In contrast to approaches that introduce low-rank branches externally, QwT [21] integrates lightweight linear layers within each decoder layer block of LLMs, directly compensate quantization errors, offering new insights for quantization error mitigation.
2.3 Symmetric vs. asymmetric
Uniform quantizers can be broadly categorized into symmetric and asymmetric schemes, depending on how the zero point is handled during the mapping from floating-point to integer values.
(a) Symmetric quantization
(b) Asymmetric quantization
Figure 3: Symmetry of quantization
Symmetric quantization forces the zero point to align with the floating point, resulting in a quantization range that is symmetric with respect to zero. The quantization process is typically defined as:
$X_{q}=\text{round}\left(\frac{X}{\Delta}\right),\quad\Delta=\frac{\max(|X|)}{2^{N}-1}$ (12)
where the floating-point tensor $X$ is scaled by a parameter $\Delta$, which is computed from the distribution of $X$. Eq. (12) presents a commonly used Min-Max method for computing the quantization step size $\Delta$, where $N$ denotes the bit-width of the quantization precision.
The dequantization process can be formulated as:
$\text{dequant}(X)=X_{q}\Delta$ (13)
$\text{dequant}(Y_{q})=X_{q}W_{q}\Delta_{X}\Delta_{W}=Y_{q}\Delta_{X}\Delta_{W}$ (14)
It is evident that symmetric quantization enables a very simple dequantization process. For linear layers, dequantization can be efficiently performed by multiplying the quantized output with the product of the scale factors for the weights and activations like Eq. (14). This simplifies computation and is often preferred in hardware implementations. However, symmetric quantization may lead to suboptimal representation when the data distribution is not centered around zero.
Asymmetric quantization allows the zero point to shift dynamically according to the minimum and maximum values of the data. This enables a tighter fit to the data distribution, especially when it is skewed, at the cost of slightly more complex computation.
Asymmetric quantization incurs additional computational cost during dequantization, while it often outperforms symmetric quantization in terms of accuracy. The dequantization process can be formulated as:
$\text{dequant}(X)=(X_{q}-Z)\Delta$ (15)
$\text{dequant}(Y_{q})=(X_{q}-Z_{X})\Delta_{X}\cdot(W_{q}-Z_{W})\Delta_{W}$
$=Y_{q}\Delta_{X}\Delta_{W}+(Z_{X}W_{q}+Z_{W}X_{q}+Z_{X}Z_{W})\Delta_{X}\Delta_{W}$ (16)
As such, the selection of quantization symmetry should be made with consideration of the specific use case and performance-complexity trade-offs.
2.4 Granularity of quantization
The granularity of quantization often largely effects the low-bit quantization performance. Commonly adopted granularities include per-tensor, per-channel or per-token, and per-group, as illustrated in Fig. 4.
Figure 4: Granularity of quantization: (1) Per-tensor quantization; (2) Per-channel quantization; (3) Per-group quantization.
In per-tensor quantization, a single set of affine transformation parameters is applied to the entire tensor, with the scale ($\Delta$) and zero point ($Z$) computed from the tensor minimum and maximum values. A finer-grained alternative is per-channel quantization for weights and per-token quantization for activations, where each input channel of the weights or each token of the activations shares the same quantization parameters. For low-bit quantization scenarios, even finer granularity, per-group quantization, is often employed to further enhance performance. In this case, the elements within a single channel or token are subdivided into groups of a predefined size, with each group quantized independently using its own parameters.
Finer granularity typically leads to improved performance but also incurs higher memory overhead due to the need to store more quantization parameters. Besides, as QServe [17] mentioned, per-group quantization will also raise computational cost in hardware level. Therefore, while finer granularity can enhance accuracy, it may also reduce the overall efficiency of the system. In weight-only quantization, papers such as GPTQ [8], AWQ [11], and QuIP [9] all employ per-group quantization to further enhance performance. For weight-activation quantization, Atom [1] also adopts a per-group quantization strategy. However, with the introduction of rotation-based methods in QuaRot [13] and SpinQuant [3], the performance of per-channel quantization has been significantly improved. Inspired by this, subsequent papers like FlatQuant [12], ResQ [19], and OSTQuant [4] have also adopted per-channel granularity for quantization.
2.5 MXFP4 data format
Recently, the NVIDIA’s GeForce RTX 50 Series based on NVIDIA’s Blackwell architecture [22] is released, which has enabled very low-precision operations using FP4 variants like MXFP4 [23, 24] and NVFP4 [25]. MXFP4 stores each value in an E2M1 format, consisting of one sign bit, one mantissa bit, and two exponent bits. Unlike the INT4 format, MXFP4 employs a fixed granularity, utilizing symmetric per-group quantization with a group size of 32. Each group is associated with an E8M0 scaling factor, and this rescaling process is hardware-integrated.
The value of an encoding MXFP4 data is represented as:
$\text{value}=\begin{cases}(-1)^{S}\times 2^{1-\text{bias}}\times(0+2^{-m}\times M)&\text{if }E=0\\ (-1)^{S}\times 2^{E-\text{bias}}\times(1+2^{-m}\times M)&\text{if }E>0\end{cases}$ (17)
where $S,E,M$ corresponds to the sign,exponent and mantissa fields, respectively. The bias is the exponent bias which is set to 1 for MXFP4. So the MXFP4 encoding values are [-6, -4, -3, -2, -1.5, -1, -0.5, 0, 0.5, 1, 1.5, 2, 3, 4, 6].
The quantization process of MXFP4 is similar to the integer quantization process. First, we compute the scaling factor, using the same way as symmetric quantization situation mentioned above that computes from the maximum value of a group of values. Then the scaling factors need to be cast to E8M0 format. After scaling, the values are quantized to the nearest MXFP4 encoding values.
MXFP4 is enable to better handle long-tail distributions compared to INT4, as its values are non-uniformly distributed. In LLMs, both weights and activations approximately follow normal distributions, with activations particularly exhibiting long-tail characteristics. Fig. 5 shows the weight distribution under BF16, INT4, and MXFP4. A clear observation is that under INT4 quantization, values with larger absolute magnitudes represent fewer data points, whereas MXFP4 more fully utilizes the representational capacity of 4-bit precision. Therefore, MXFP4 is more suitable than INT4 as a quantization data type.
(a) Full-precision weight
(b) INT4 quant weight
(c) MXFP4 quant weight
Figure 5: Weight distribution of a q_proj layer of LlaMA-3.2-1B. (a): Weight distribution in full-precision(BF16). (b): Weight distribution after INT4 quantization (group size=32), compared with MXFP4. (c): Weight distribution after MXFP4 quantization.
LLM-FP4 [26] studies FP4 precision quantization for LLMs, including searching for the optimal exponent bias and maximal quantization value. It also proposed a pre-shifted exponent bias technique to address the issue of high inter-channel variance. FP4-Quantization [27] improves the FP4 data representation while introducing a method called LREC (low-rank error calibration) to quantize weights into FP4.
3 Extensive Evaluations
Quantization settings We conduct experiments using W4A4 precision, comparing the results with the origin precision (BF16 or FP16). In the main results, we use per-token or per-group (group size=128) asymmetric quantization for activations. For weight quantization, we use per-channel or per-group (group size=128) symmetric quantization, where we extract the clipping ratio using a linear search over the squared error.
Models and tasks. We conduct experiments on LLaMA-3.2-1B, LLaMA-3.2-3B, and Qwen-2.5-3B. We evaluate their perplexity on the WikiText-2 dataset [28] and assess their accuracy on several zero-shot evaluation tasks, including commonsense reasoning, mathematics, sentiment analysis, and question answering. Specifically, we utilize the following datasets for evaluation: MMLU [29], SST2 [30], MathQA [31], ARC‑C [32], HellaSwag [33], WinoGrande [34], BoolQ [35], LAMBADA [36], and PIQA [37].
Implementation. To calibrate the GPTQ [8] quantization and the low-rank compensation branch, we use 128 samples from WikiText-2 with a sequence length of 2048. The training of the learnable orthogonal and scaling matrices follows OSTQuant [4] procedure, using 1,000 samples from WikiText-2 and optimized with RiemannAdam. For the low-rank branch, we set the rank k=32 and store the low-rank matrices in FP16. For the non-trainable scaling matrix, we follow the SmoothQuant [6] calibration scheme with $\alpha$=0.5.
3.1 Overall performance
Evaluate Methods Channel Group-128
DN EC Wiki2‑PPL($\downarrow$) 0‑shot9 Avg.($\uparrow$) Wiki2‑PPL($\downarrow$) 0‑shot9 Avg.($\uparrow$)
BF16 9.76 54.78 9.76 54.78
rotation - 17.83 45.24 14.84 47.51
rotation low-rank 17.16 45.47 14.5 47.04
rotation GPTQ 12.95 47.79 11.65 49.38
rotation GPTQ+low-rank 12.81 48.35 11.61 50.49
rotation+scaling - 14.88 46.41 13.3 49.22
rotation+scaling low-rank 14.58 46.76 13.07 49.95
rotation+scaling GPTQ 11.8 48.89 10.89 50.63
rotation+scaling GPTQ+low-rank 11.73 49.06 10.87 51.83
Table 3: Evaluation of Quantization Method Combinations, in which rotation or scaling is optimized, in terms of Perplexity on WikiText2 and averaged accuracy on nine Zero-Shot tasks across per-channel and per-group settings of LlaMA-3.2-1B.
Evaluate Methods Channel Group-128
DN EC Wiki2‑PPL($\downarrow$) 0‑shot9 Avg.($\uparrow$) Wiki2‑PPL($\downarrow$) 0‑shot9 Avg.($\uparrow$)
BF16 9.76 54.78 9.76 54.78
rotation 20.31 43.26 15.18 46.62
rotation low‑rank 19.84 43.09 14.83 46.87
rotation GPTQ 13.31 48.02 11.96 50.05
rotation GPTQ+low‑rank 13.28 48.37 11.98 50.48
rotation+scaling 20.96 42.87 16.25 45.01
rotation+scaling low‑rank 21.97 43.07 15.55 45.46
rotation+scaling GPTQ 14.04 47.53 12.23 48.95
rotation+scaling GPTQ+low‑rank 13.91 46.46 12.17 49.41
Table 4: Evaluation of Quantization Method Combinations, in which rotation or scaling is non-optimized, in terms of Perplexity on WikiText2 and averaged accuracy on nine Zero-Shot tasks across per-channel and per-group settings of LlaMA-3.2-1B.
We select two of the most widely-used pre-quantization transformation techniques, scaling and rotation, and combined them with two quantization error mitigation methods: GPTQ and low-rank branch compensation. A series of experiments are conducted to evaluate these combinations. Table 3 presents the results of using optimized scaling and rotation, illustrating the performance of different method combinations. Table 4 presents the results obtained with randomly initialized rotations and calibrated scaling factors.
Our experiments reveal that the best performance is achieved when both optimized scaling and rotation are applied in conjunction with GPTQ and the low-rank branch. This configuration yields a perplexity of 11.73 in per-channel quantization for Llama-3.2-1B model, which is close to the BF16 baseline of 9.76.
Between the two quantization error mitigation methods, GPTQ clearly plays a more significant role than low-rank adaptation. Under the optimal pre-quantization transformation setting (optimized rotation and scaling), using GPTQ alone results in perplexity of 11.8, whereas using only the low-rank method yields 14.58, under the per-channel settings. When both methods are used together, performance improves further, from 11.8 to 11.73 in the per-channel setting, and from 10.89 to 10.87 in the per-group setting, narrowing the gap to the full-precision model.
In addition, using both optimized rotation and scaling improves performance over using only optimized rotation. However, when employing non-optimized rotation and scaling, rotation alone tends to yield better results than the combination, which indicates that calibrated scaling may not performs well together with rotation in W4A4 settings.
Model precision Wiki2 PPL($\downarrow$) Zero-Shot Accuracy($\uparrow$)
General Sentiment Math Reasoning QA Avg.
MMLU SST2 MathQA PiQA HellaS WinoG BoolQ Lambada Arc-C
Llama- 3.2-1B fp16 9.76 36.76 66.86 28.94 74.65 63.67 60.38 63.67 62.1 36.01 54.78
w4a4 11.73 29.21 55.28 25.43 69.15 57.93 56.43 59.02 56.78 32.34 49.06
Llama- 3.2-3B fp16 7.82 54.06 73.62 34.47 77.48 73.58 69.46 72.97 70.08 46.08 63.53
w4a4 8.9 46.7 68.58 30.55 73.99 69.54 65.43 68.75 65.32 41.55 58.93
Qwen- 2.5-3B fp16 8.03 65.13 90.37 37.15 78.73 73.55 68.75 77.58 66.8 47.35 67.27
w4a4 9.22 56.32 69.15 32.5 74.16 67.52 62.12 65.41 57.77 42.83 58.64
Table 5: Comparision of s ymmetric and asymmetric quantization in Terms of Perplexity on WikiText2 and averaged accuracy on Zero-Shot tasks of LlaMA-3.2-1B
In Table 5, we further conduct experiments on different models: LLaMA-3.2-1B, LLaMA-3.2-3B, and Qwen-2.5-3B, using the optimized scaling and rotation, and combined them with GPTQ and an additional low-rank branch.
3.2 Symmetric vs. asymmetric
Granularity Metric rotation+scaling+GPTQ rotation+scaling+GPTQ+low-rank
Sym W-Asym A-Asym Asym Sym W-Asym A-Asym Asym
channel Wiki2-PPL($\downarrow$) 12.73 12.7 11.8 11.7 12.63 12.65 11.73 11.64
Arc-Challenge Score(%) 31 31.7 33.3 33.7 31.9 29.9 32.3 32.7
Table 6: Comparision of symmetric and asymmetric quantization in terms of Perplexity on WikiText2 and Arc [32] accuracy of LlaMA-3.2-1B
In Table 6, we evaluate the performance of applying symmetric or asymmetric quantization to weights and activations, respectively. The results demonstrate that asymmetric quantization for activations leads to a significant improvement over symmetric quantization, while the performance gain for weights is relatively minor. We attribute this to the fact that activation distributions tend to be asymmetric, whereas weights are generally more symmetric in distribution.
Therefore, an effective configuration is to use symmetric quantization for weights and asymmetric quantization for activations. This setup not only achieves much better performance than fully symmetric quantization, but also reduces the computational overhead associated with dequantization process, compared to fully asymmetric quantization for both weights and activations.
3.3 Granularity
As illustrated in Fig. 6, we conduct a comprehensive comparison of per-group quantization performance across varying group sizes, measured by perplexity on WikiText2. The experimental results reveal that finer quantization granularity yields superior model performance (lower perplexity), albeit at the expense of elevated storage requirements for quantization parameters. This trade-off stems primarily from the FP16-format scaling factors, whose storage overhead scales inversely with group size. Quantitative analysis of these additional bit costs is presented in Table 7, which compares the average bit overhead introduced by scaling factors under symmetric weight quantization.
Figure 6: Comparision of different group sizes of quantization in terms of Perplexity on WikiText2 and averaged accuracy on Zero-Shot tasks of LlaMA-3.2-1B
Granularity group-32 group-64 group-128 group-256 group-512
Average Extra bits 0.5 0.25 0.125 0.0625 0.03125
Table 7: Average extra bits introduced by scaling factors of different granularities
3.4 MXFP4 vs. INT4
Extending experiments to the MXFP4 format, we follow MXFP4 configurations implementing symmetric quantization with group size 32. For activations, we use a clipping threshold of 3/4. For weights, we follow the experiments of INT4 quantization, using a linear search over the squared error to find a best clipping ratio. First, we apply no extra pre-quantization transformation or quantization error mitigation process, conducting experiments using only RTN (round to nearest). Table 8 demonstrates that, MXFP4 shows much better performance than INT4 (per-channel).
Error compensation Data Format
INT4 (channel) INT4 (g-32) MXFP4
- 923.72 14.19 15.19
GPTQ 1007.33 14.73 13.18
low-rank 723.54 14.04 15.59
GPTQ+low-rank 578.61 12.48 12.72
Table 8: Evaluation of MXFP4 format quantization without using pre-quantization transformation, in terms of Perplexity on WikiText2 of LlaMA-3.2-1B.
Evaluate Methods WikiText2PPL($\downarrow$)
Data Normalization Error Compensation INT4(channel) MXFP4
scaling GPTQ 134.85 13.06
scaling low-rank 165.4 14.83
scaling GPTQ+low-rank 124.56 12.64
optimized rotation+scaling GPTQ 11.8 12.61
optimized rotation+scaling low-rank 14.58 15.53
optimized rotation+scaling GPTQ+low-rank 11.73 12.29
Table 9: Evaluation of MXFP4 format quantization, in terms of Perplexity on WikiText2 of LlaMA-3.2-1B.
Further, we conduct experiments using both pre-quantization transformation and quantization error mitigation techniques. Results in Table 9 demonstrate that both pre-quantization transformation and quantization error mitigation techniques achieve competitive performance under MXFP4 quantization, confirming the direct applicability of INT4 quantization methods to FP4. Crucially, scaling-based approaches proved more effective than rotation-based methods for MXFP4 –- a reversal of their relative efficacy observed in INT4 quantization scenarios.
4 Conclusions
In this paper, we have presented an experimental evaluation of Large language model quantization. We systematically decouple the commonly used quantization methods into two key steps: pre-quantization transformation and quantization error mitigation. After a comprehensive review of existing approaches, we choose scaling and rotation in pre-quantization transformation step, GPTQ and low-rank compensation in quantization error mitigation step to further conduct experimental evaluations on the same ground. We found that uses the optimized rotation and scaling, together with two compensation methods demonstrating the best performance. For quantization settings, asymmetric quantization and smaller group sizes yield better results. For MXFP4 quantization, we can adopt methods used in INT4 quantization. However, using scaling-based pre-quantization transformation demonstrates superior performance compared to INT4. To achieve optimal performance under MXFP4 quantization, further investigation is required to explore its full potential and limitations.

### v2

@@
 Though many papers have reported breakthrough performance, they may not conduct experiments on the same ground since one quantization method solution usually contains multiple components.
-In addition, it is also important to analyze the theoretical connections among numerous existing methods to gain in-depth understanding in this field.
+In addition, it is also important to analyze the connections among numerous existing methods to gain in-depth understanding of this field.
 To bridge these gaps, in this paper, we timely conduct an extensive review of the state-of-the-art methods and also perform comprehensive evaluations on the same ground to ensure fair comparisons of these methods.
 To our knowledge, this fair and extensive investigation remains critically important yet significantly underexplored in the research community.
-To better understand the theoretical connections, we decouple the published quantization methods into two steps: pre-quantization transformation and quantization error mitigation.
+To better understand the connections, first, we decouple the published quantization methods into two steps: pre-quantization transformation and quantization error mitigation.
 We define the former as a preprocessing step applied before quantization to reduce the impact of outliers, making the data distribution flatter and more suitable for quantization.
 Quantization error mitigation, on the other hand, involves techniques that offset the errors introduced during quantization, thereby enhancing model performance.
-Meanwhile, we extensively evaluate and analyze the impact of different components of quantization methods.
-Additionally, we also timely analyze and evaluate the latest MXFP4 data format and its performance.
-Our experimental results demonstrate that the optimized rotation and scaling yield the best performance for pre-quantization transformation, and combining low-rank compensation with GPTQ occasionally outperforms using GPTQ alone for quantization error mitigation.
-Furthermore, we explore the potential of the latest MXFP4 quantization and reveal that the optimal pre-quantization transformation strategy for INT4 does not generalize well to MXFP4, inspiring further investigation in this field.
+Second, we extensively evaluate and analyze the impact of different settings of quantization methods, including granularity and symmetry.
+Third, we also timely analyze and evaluate the latest MXFP4 and NVFP4 data format and its performance.
+Our experimental results first demonstrate that the optimized rotation and scaling yield the best performance for pre-quantization transformation, and combining low-rank compensation with GPTQ occasionally outperforms using GPTQ alone for quantization error mitigation.
+Second, as the granularity becomes finer, the performance improves, but this also incurs additional storage overhead.
+We also verified that applying asymmetric quantization to weights yields much smaller benefits compared to applying it to activations, which is consistent with the common practice of using asymmetric quantization primarily to activations.
+Third, we explore the potential of the latest FP4 quantization and find that the format and precision of scaling factors have a significant impact on the performance.
+We also show the performance of the optimal pre-quantization transformation strategy for INT4 on FP4, revealing that rotation-based methods demonstrate little improvement in performance for MXFP4 and NVFP4, inspiring further investigation in this field.
 keywords:
@@
 Representative settings include W4A4 [1, 2, 3, 4], W8A8 [5, 6, 7] and W4A16 [8, 9, 10, 11]: W4A4 quantizes both weights and activations to 4-bit precision (INT4); W8A8 and W4A16 quantize weights to INT8 and INT4 and activations to INT8 and FP16/BF16 respectively.
-Among these 3 settings, W4A4 is the current research focus because it provides the best performance in terms of both memory efficiency and inference speed, with the greatest research significance.
-However, despite the wide range of methods explored in previous studies, W4A4 quantization still struggles to achieve lossless performance, especially for smaller LLMs, such as those with 1B, 3B, and 7B parameters, which often results in increased perplexity and degraded performance on various tasks.
-Furthermore, LLaMA3 experiences a larger performance drop after quantization compared to LLaMA2 when using many popular quantization methods [3, 4, 12], facing a greater demand for better W4A4 quantization methods.
-Therefore, W4A4 quantization for LLMs still remains a challenging task.
-To review and evaluate the state-of-the-art quantization methods, in particular, under the setting of W4A4, we find two main characteristics of these methods.
-First, existing methods are often combinations of multiple techniques, since using a single technique alone typically fails to achieve excellent quantization performance.
-In SpinQuant [3], after applying trained rotations to adjust the model, the GPTQ [8], a gradient-based quantization error mitigation method is further applied to quantize the rotated weights.
-Similarly, OSTQuant [8] employs multiple techniques to adjust the weights and activations, including rotation and scaling.
-At the same time, there are many weight-only quantization methods, some of which can be combined with pre-quantization transformation techniques and extended to weight-activation quantization.
-AWQ [11] shares similar pre-quantization transformation (scaling) with SmoothQuant [6], even though their motivations differ.
-Also, QuIP# [10] shares similar pre-quantization transformation (random hadamard rotation) with QuaRot [13].
-Apart from the methods that involve both weight and activation transformations, there are also approaches that are orthogonal to pre-quantization transformation that can be used to compensate for the loss between quantized models and full precision models, such as GPTQ [8] and low-rank branches computed from quantization error [14, 15, 16].
-Second, quantization involves a variety of configurations, such as granularity and symmetric or not.
+Among these three settings, W4A4 is currently the main focus of research, as it holds great potential for improving memory efficiency and inference speed.
+However, applying lossless W4A4 quantization to LLMs remains a challenging problem.
+Thus, we concentrate our efforts on exploring W4A4 quantization.
+In this paper, we mainly investigate three dimensions.
+First, we find that existing methods are often combinations of multiple techniques, since using a single technique alone typically fails to achieve excellent quantization performance.
+OSTQuant [8] and FlatQuant [12] employ multiple techniques to adjust the weights and activations, including rotation and scaling, and also apply the GPTQ [8], a gradient-based method for mitigating quantization error, to quantize the transformed weights.
+Besides, many techniques share similar transformations, for instance, Quip# [10] and QuaRot [13] both use rotation, also, AWQ [11] and SmoothQuant [6] both use scaling.
+To achieve a better understanding of existing methods and make our evaluation on the same ground, we decompose the quantization process into two key steps: pre-quantization transformation and quantization error mitigation.
+The pre-quantization transformation step refers to preprocessing the data before quantization in order to mitigate the effect of outliers, making the data distribution smoother and more quantization-friendly.
+The quantization error mitigation step involves applying techniques to compensate for the errors introduced by quantization, thereby reducing the overall quantization error and improving performance.
+For each step, we select several state-of-the-art techniques and conduct an evaluation of the effectiveness of their combinations.
+Second, quantization involves a variety of configurations, such as granularity and whether symmetric or not.
 Different settings can lead to different performance outcomes and often involve trade-offs between efficiency and accuracy.
 For different precision levels, the optimal quantization settings may vary.
-Evaluating the performance under various quantization configurations for W4A4 quantization is of great importance.
-In this paper, as aforementioned in the first characteristic, the existing quantization solutions are very complicated and they usually consists of many algorithms.
-To make our evaluation on the same ground and achieve a better understanding of existing methods, we decompose the quantization process into two key steps: pre-quantization transformation and quantization error mitigation.
-The pre-quantization transformation step refers to preprocessing the data before quantization in order to mitigate the effect of outliers making the data distribution smoother and more quantization-friendly.
-The quantization error mitigation step involves applying techniques to compensate for the errors introduced by quantization, thereby reducing the overall quantization error and improving performance.
-For each step, we select several state-of-the-art techniques and conduct an evaluation of the effectiveness of their combinations.
-Second, existing solutions usually do not use exactly the same setting, e.g.
-granularity, leading to unfair comparison.
+And existing solutions usually do not use exactly the same setting, leading to unfair comparison.
 To bridge this gap, we conduct a comprehensive evaluation on the same quantization configurations and provide recommendations on selecting appropriate configurations.
-And we integrate the methods used in this paper in our code library to facilitate performance evaluation.
-Furthermore, the NVIDIA’s GeForce RTX 50 Series latestly released based on the new RTX Blackwell architecture includes support for the FP4 data format, including NVFP4 and MXFP4, in its new Tensor Cores, which show great advantages over INT4.
+And we integrate the methods used in this paper into our code library to facilitate performance evaluation.
+Third, NVIDIA’s GeForce RTX 50 Series, which is newly released and based on the Blackwell architecture, includes support for the FP4 data format (e.g., NVFP4 and MXFP4) in its Tensor Cores, showing significant advantages over INT4.
 However, the research on FP4 has not been well explored.
 To bridge this gap, we conduct research on weight-activation FP4 quantization using the popular methods in INT4, showing the great timeliness of our research.
-Methods Data Normalization Error Compensation
-SmoothQuant [6] scale -
-Outlier Suppression+ [7] scaling + shifting -
+For FP4 quantization, we also observe that the format and precision of scaling factors have a significant impact on the performance.
+2 Review on Existing Quantization Methods
+In this section, we first review recent quantization algorithms in Section 2.1 in terms of two steps: pre-quantization transformation and quantization error mitigation.
+Table 1 shows our decomposition of the popular quantization methods.
+Section 2.2 focuses on settings, including symmetry and granularity of quantization methods.
+Meanwhile, Table 2 summarizes the settings of existing methods, showing that they are not compared on the same ground.
+Section 2.3 discusses FP4 quantization and its variants, including MXFP4 and NVFP4.
+Methods Pre-quantization Transformation Quantization Error Mitigation
+SmoothQuant [6] Scaling RTN
+Outlier Suppression+ [7] Shifting + Scaling RTN
 GPTQ [8] - GPTQ
@@
 QuIP [9] Rotation GPTQ
-AWQ [11] scaling -
+AWQ [11] Scaling RTN
 Atom [1] Reorder GPTQ
-OmniQuant [2] optimized scaling + shifting GPTQ
+OmniQuant [2] Shifting+Scaling GPTQ
 Qserve [15] Scaling + Rotation GPTQ
-LQER [18] - low-rank
+LQER [16] Scaling Low-rank
 QuaRot [13] Rotation GPTQ
-SpinQuant [3] optimized rotation GPTQ
+SpinQuant [3] Rotation GPTQ
 ResQ [17] Reorder + Rotation GPTQ
-FlatQuant [12] optimized scaling + rotation GPTQ
-OSTQuant [4] optimized scaling + rotation GPTQ
+FlatQuant [12] Scaling + Rotation GPTQ
+OSTQuant [4] Scaling + Rotation GPTQ
 Table 1: Two steps decomposition of quantization methods.
-2 Review on Existing Quantization Methods
-In this section, we first review recent quantization algorithms in Section 2.1 and Section 2.2 in terms of two steps: pre-quantization transformation and quantization error mitigation.
-Table 1 shows our decomposition of the popular quantization methods.
-Section 2.3 and Section 2.4 focus on settings, including symmetricity and granularity of quantization methods.
-Meanwhile, Table 2 summarizes the quantization settings reported in the experimental results of these methods.
+2.1 Two-steps of quantization
+2.1.1 Pre-quantization transformation
+Pre-quantization transformation is a step before quantization, applying operations such as shifting, scaling, and rotation to process weights and activations.
+These transformations aim to flatten and smooth the hard-to-quantize tensors, thereby enhancing the quantization performance.
+LLMs are difficult to quantize due to the outliers in the activations.
+Compared to most of the activation values, the magnitude of outliers is substantially higher, often exceeding 100x larger in some layers.
+The outliers will dominate the maximum magnitude measurement, leading to low effective quantization levels.
+This pre-quantization transformation step is illustrated in Fig.
+1.
 Methods Precision Symmetric Granularity
@@
 OSTQuant [4] W4A4 T F C
-Table 2: Symmetry and granularity of quantization methods. In granularity, T stands for per-tensor, C represents per-channel or per-token, and G means per-group
-2.1 Pre-quantization transformation
-Pre-quantization transformation is a step before quantization, applying operations such as shifting, scaling, and rotation to process weights and activations.
-These transformations aim to flatten and smooth the hard-to-quantize tensors, thereby enhancing the quantization performance.
-LLMs are difficult to quantize due to the outliers in the activations.
-Compared to most of the activation values, the magnitude of outliers is substantially higher, often exceeding 100x larger in some layers.
-The outliers will dominate the maximum magnitude measurement, leading to low effective quantization levels.
-This pre-quantization transformation step is illustrated in Fig.
-1.
-Figure 1: Pre-quantization transformation: Transforming activations and weights to render them more easier to quant
-2.1.1 Shifting
-Activations distribution is often asymmetric, and this asymmetry varies across different channels.
+Table 2: Symmetry and granularity of quantization methods. In granularity, T stands for per-tensor, C represents per-channel or per-token, and G means per-group.
+Figure 1: Pre-quantization transformation: a process that transforms activations and weights to make them easier to quantize.
+The left graph illustrates a tensor that is difficult to quantize due to the presence of numerous outliers, while the right graph shows that after the pre-quantization transformation, the tensor becomes more uniformly distributed, making quantization easier.
+Shifting
+Activation distribution is often asymmetric, and this asymmetry varies across different channels.
 Aligning the center of each channel can improve the performance of quantization, since we quantize activations per-token rather than along the channel dimension.
 Outlier Suppression+ [7] proposes shifting activation values to eliminate the asymmetry feature across channels.
-The shifting process is to use a vector that transfer the channels of the activation.
+The shifting process is to use a vector $T$ that transfers the channels of the activation.
 For a linear layer, it is formulated as:
-$Y=XW+B=(\underbrace{(X-T)}_{\hat{X}})W+(\underbrace{TW+B}_{\hat{B}})$ (1)
+[…]
 where the layer input $X\in\mathbb{R}^{T\times C_{in}}$ is shifting and becomes $\hat{X}$.
 $T\in\mathbb{R}^{1\times C_{in}}$ is a vector representing the channel-wise shifting parameters.
-To keep mathematical equivalent, an additional term, the product of $T$ and $W\in\mathbb{R}^{C_{in}\times C_{out}}$, is added to the bias $B\in\mathbb{R}^{1\times C_{out}}$.
-Outlier Suppression+ [7] uses calibration to get the vector $T$, which is computed channel-wise as the average of its minimum and maximum values:
-$t_{i}=\frac{\min(X_{i})+\max(X_{i})}{2}$ (2)
-where $t_{j}$ denotes the $j$-th element of $T$, corresponding to the $j$-th channel.
-$X_{j}$ represents the $j$-th input channel of $X$.
-OmniQuant [2] employs learnable channel-wise shifting for outlier suppression, combined with the learnable scaling technique discussed in Section 2.1.2, and jointly optimize these two kinds of pre-quantization transformation parameters.
-2.1.2 Scaling
+To keep mathematical equivalence, an additional term, the product of $T$ and $W\in\mathbb{R}^{C_{in}\times C_{out}}$, is added to the bias $B\in\mathbb{R}^{1\times C_{out}}$.
+Methods Shifting Strategy Shifting parameters $T$
+OS+ [7] Pre-scaling shifting $T_{i}=\frac{\min(X_{i})+\max(X_{i})}{2}$
+OmniQuant [2] Pre-scaling shifting Gradient-based optimization
+Table 3: Comparison of shifting-based quantization methods.
+As shown in Table 3, Outlier Suppression+ [7] uses calibration to get the vector $T$, while OmniQuant [2] employs learnable channel-wise shifting for outlier suppression, combined with the learnable scaling technique discussed in the next paragraph, and jointly optimizes these two kinds of pre-quantization transformation parameters.
+Scaling
 Scaling is a commonly used pre-quantization transformation method, which scales activations in different input channels before quantization, and inversely scales the corresponding weight channels to make the results stay unchanged.
-SmoothQuant [6] first propose using scaling to smooth the quantization difficulty from activations to weights.
+SmoothQuant [6] first proposes using scaling to smooth the quantization difficulty from activations to weights.
 Compared to weights, activations have more outliers that make quantization harder.
@@
 $Y=XW=(\underbrace{X\cdot S^{-1}}_{\hat{X}})(\underbrace{S\cdot W}_{\hat{W}})=\hat{X}\hat{W}$ (2)
-where the weights of a linear layer $W$ and the layer input $X$ is scaling and becomes $\hat{X}$ and $\hat{W}$.
-$S$ is a diagonal matrix composed by scaling factors of different channels, which can be obtained in various ways.
-SmoothQuant [6] uses calibration to compute $S$.
-The calibration process can be formulated as:
-$s_{j}=max(|X_{j}|)^{\alpha}/max(|W_{j}|)^{1-\alpha}$ (4)
-where $X_{j}$ and $W_{j}$ denote the $j$-th input channel of $X$ and $W$, respectively.
-$s_{j}$ is the $j$-th diagonal element of $S$, corresponding to the $j$-th channel.
-$\alpha$ is a hyperparameter that controls the degree to which outliers are transferred between activations and weights.
-AWQ [11], as a weight-only quantization method, approaches the problem from the perspective of protecting salient weights, applying activation-aware scaling.
-Unlike adopting the hand-crafting parameters in calibration, the scaling operation in OmniQuant [2] uses learnable scaling, where $S$ is learnable parameters that are optimized with block-wise quantization error minimization.
+where $S$ is a diagonal matrix composed of scaling factors of different channels, which can be obtained in various ways.
+Methods Scaling Strategy Scaling parameters $S$
+SmoothQuant [6] Scaling only $\begin{aligned} &S_{i}=\frac{\max(|X_{i}|)^{\alpha}}{\max(|W_{i}|)^{1-\alpha}}\end{aligned}$
+AWQ [11] Scaling only $\begin{aligned} &S_{i}=|X_{i}|_{avg}^{\alpha}\\ &\alpha^{*}=\arg\min_{\alpha}\mathcal{L}(S_{Xi}^{\alpha})\end{aligned}$
+OS+ [7] Post-shifting scaling $\begin{aligned} &S_{i}=\max\left(1.0,\frac{\max(X_{i}-T_{i})}{t}\right)\\ &t^{*}=\arg\min_{t}\mathcal{L}(S_{i})\end{aligned}$
+OmniQuant [2] Post-shifting scaling Gradient-based optimization
+FlatQuant [12] Pre-rotation scaling Calibration initialization and gradient-based optimization
+OSTQuant [4] Post-rotation scaling Gradient-based optimization
+Table 4: Comparison of scaling-based quantization methods.
+As shown in Table 4, SmoothQuant [6] uses calibration to compute $S$.
+AWQ [11], as a weight-only quantization, approaches the problem from the perspective of protecting salient weights.
+Its scaling parameter is determined by the average magnitude of the activation, while the optimal exponent $\alpha^{*}$ is found by minimizing the loss through grid searching.
+Outlier Suppression+ [7] also uses grid searching optimization to find the optimal scaling after applying shifting.
+Unlike adopting the hand-crafting parameters in calibration, Outlier Suppression+ [7] uses grid searching optimization to find the optimal scaling after applying shifting.
+OmniQuant [2] optimizes the quantization parameters via gradient-based methods to minimize the block-wise quantization error.
 FlatQuant [12] and OSTQuant [4] both use learnable per-channel scaling, but at different stages of pre-quantization transformation, leading to different practical meanings.
-Both perform a diagonal scaling operation, yet FlatQuant scales before rotation, allowing its scaling factors to be initialized via Eq.
-(4), functionally similar to SmoothQuant.
-In contrast, OSTQuant applies scaling after rotation, meaning this diagonal scaling is a post-rotation adjustment rather than smoothing the original channels.
-2.1.3 Rotation
-The rotation operation is a technique that adjusts data distribution by multiplying the data matrix with an orthogonal matrix.
-This operation largely changes the activation distributions, reduces outliers in the data and thereby enhancing quantization accuracy.
+Rotation
+The rotation operation is a technique that adjusts data distribution by multiplying the data matrix by an orthogonal matrix.
+This operation largely changes the activation distributions, reduces outliers in the data, and thereby enhances quantization accuracy.
 The rotation process in a linear layer can be formulated as:
@@
 To preserve the model’s output, the weight $W$ is simultaneously transformed by $O^{\top}$, which is the transpose of $O$.
-QuIP [9] first introduced rotation with random orthogonal matrices to reduce incoherence in weight-only quantization.
+Methods Rotation Strategy Rotation parameters $O$
+QuIP [9] Rotation only Random orthogonal matrices
+QuIP# [10] Rotation only Random Hadamard matrices
+QuaRot [13] Rotation only Random Hadamard matrices
+SpinQuant [3] Rotation only Gradient-based optimization
+FlatQuant [12] Post-scaling Rotation Gradient-based optimization
+OSTQuant [4] Pre-scaling Rotation Gradient-based optimization
+Table 5: Comparison of rotation-based quantization methods.
+As shown in Table 5, QuIP [9] first introduced rotation with random orthogonal matrices to reduce incoherence in weight-only quantization.
 QuIP# [10] replaces this with random Hadamard matrices for greater efficiency.
@@
 SpinQuant [3] further improves stability by optimizing the rotation matrix on the Stiefel manifold to minimize quantization-induced performance variance, while still defaulting to random Hadamard for a few online rotations.
-Rotation is a highly effective pre-quantization transformation technique for reducing outliers in low-bit quantization, and it has also been widely adopted in many other quantization methods (e.g., ResQ [19], QServe [17], FlatQuant [12] and OSTQuant [4], etc), which combine rotation with other pre-quantization transformation techniques to further improve performance.
-2.2 Quantization error mitigation
-The quantization error mitigation step employs techniques that mitigate the error introduced by quantization by compensating them.
-This error between the original model and the quantized model can be computed from calibration datasets, and be compensated to the original weights in linear layers or through low-rank branches.
+Rotation is a highly effective pre-quantization transformation technique for reducing outliers in low-bit quantization, and it has also been widely adopted in many other quantization methods together with other pre-quantization transformation techniques to further improve performance (e.g., ResQ [17], QServe [15], FlatQuant [12], and OSTQuant [4], etc).
+2.1.2 Quantization error mitigation
+The quantization error mitigation step employs techniques that mitigate the error introduced by quantization by compensating for it.
+This error between the original model and the quantized model can be computed from calibration datasets, and compensated to the original weights in linear layers or through low-rank branches.
 Fig.
@@
 Figure 2: Quantization error mitigation: compensate the error produced by quantization.
-(1) Self-compensation [8, 9] calculates weight compensation terms through quantization-induced loss errors; (2) Low-rank branch compensation [14, 16, 18, 15] constructs auxiliary compensation pathways via matrix factorization of post-quantization weight errors.
-2.2.1 Self compensation
-GPTQ [8] is a quantization method for weight quantization that uses Hessian-based quantization error mitigation.
-GPTQ builds on the OBQ [20] method that solves the error introduced by quantization and the optimal weight perturbation.
-For weight quantization, the loss error introduced by quantization can be defined as:
-$\epsilon=\mathcal{L}(W_{q})-\mathcal{L}(W)=\mathcal{L}(W+\delta)-\mathcal{W}$ (6)
-where $L(W)$ is the loss function of the model evaluated at weights $W$, and $\delta$ represents the weight perturbation.
-Using a Taylor expansion, we can get the Hessian-based approximation of the quantization loss, which can be formulated as:
-$\epsilon=\nabla\mathcal{L}(W)^{\top}\delta+\frac{1}{2}\delta^{\top}H(W)\delta+\mathcal{O}(\left\|\delta\right\|^{3})\approx\frac{1}{2}\delta^{\top}H(W)\delta$ (7)
-Unlike OBQ that uses greedy order to quantize weights, GPTQ quantized all rows of weights in a fixed order.
-The GPTQ algorithm performs quantization column by column in blocks of size B.
-For each column within the block, the quantization error is computed when quantized, and then compensates on the remaining weights.
-Once all columns of a block is quantized, the remaining weights in the matrix are also updated.
-QuIP [9] also minimizes a quadratic proxy of the weight error to implement an optimal adaptive rounding procedure called LDLQ.
-In this method, an LDL decomposition of the Hessian matrix is performed to construct a feedback matrix $U$, which is then used to form a compensation term that corrects for the quantization-induced error.
-In QuIP, it is proven that for Eq.
-(7), performing an LDL decomposition on the Hessian matrix yields an optimal choice of $U$, and GPTQ is a special case of LDLQ.
-2.2.2 Low-rank branch compensation
+Here, $dq(W_{q})$ denotes dequantizing $W_{q}$ back to the original precision value space, for computational purposes.
+Self-compensation
+We define self-compensation as directly compensating for errors on the original weights.
+Existing methods that employ self-compensation are generally based on analyzing the loss error introduced by quantization and make use of its Hessian-based quadratic approximation, which can be formulated as:
+$\Delta\mathcal{L}=\mathcal{L}(dq(W_{q}))-\mathcal{L}(W)\approx\frac{1}{2}(dq(W_{q})-W)^{\top}H(dq(W_{q})-W)$ (4)
+where $\mathcal{L}(W)$ is the loss function of the model evaluated at weights $W$.
+$W_{q}$ and $dq(W_{q})$ denote the quantized weights and their dequantized version, respectively.
+Self-compensation methods need to find an optimal weight perturbation $\delta$ to add to $W$, such that the quantization loss error in Eq.
+(4) is minimized.
+Table 6 summarizes the computation strategies of $\delta$ adopted by various self-compensation quantization methods.
+Methods Weight perturbation $\delta$ computation
+OBQ [18] Layer-wise MSE
+GPTQ [8] Layer-wise MSE
+QuIP [9] LDL decomposition
+GPTAQ [19] Consider input deviation
+GQuant [20] Add gradient information
+Table 6: Comparison of self-compensation quantization methods.
+Some methods simplify the quantization problem to minimize the squared error between the quantized and original output of each layer [18, 8].
+Based on the squared error loss function, OBQ [18] and GPTQ [8] compute the Hessian matrix used in Eq.
+(4) and further calculate the optimal weight perturbation $\delta$.
+Unlike OBQ, which uses greedy orders to quantize weights, GPTQ quantizes all rows of weights in a fixed order to achieve a faster quantization process.
+Some other methods take into account additional factors beyond the linear layer-wise formulation.
+While also aiming to minimize the quadratic proxy of the weight error in Eq.
+(4), QuIP [9] performs an LDL decomposition of the Hessian matrix to construct a compensation term that corrects for quantization-induced errors.
+GPTAQ [19], built upon GPTQ, addresses the limitation that GPTQ’s layer-wise optimization objective ignores the deviation between a layer’s actual input and that under the full-precision model.
+GQuant [20] further observes that the layer-wise MSE treats all hidden features equally, and therefore incorporates per-feature gradients into the objective function to improve quantization accuracy.
+Low-rank decomposition compensation
 Beyond compensating errors directly on the original weights, an alternative approach is to introduce an auxiliary branch dedicated to error correction.
-Some prior works leverage a low-rank term to reconstruct quantization error (e.g., ZeroQuant-v2 [14], CALDERA [16], LQER [18], QERA [15], etc).
+Some prior works leverage a low-rank term to reconstruct quantization error (e.g., ZeroQuant-v2 [14], CALDERA [21], LQER [16], QERA [22], etc), where the low-rank branch can either be quantized to higher precision, such as INT8, or remain in FP16.
 After incorporating a low-rank branch, the output $Y$ is computed as:
-$Y=XW=XW_{q}+X\Delta W\approx XW_{q}+XAB$ (8)
-where $\Delta W=W-W_{q}$ denotes the weight reconstruction error introduced by quantization.
-$A\in\mathbb{R}^{C_{in}\times r}$, $B\in\mathbb{R}^{r\times C_{out}}$ are learnable low-rank matrices, and $r\ll\min(d,k)$ is the bottleneck rank.
+$Y=Xdq(W_{q})+XAB$ (5)
+where $A\in\mathbb{R}^{C_{in}\times r}$, $B\in\mathbb{R}^{r\times C_{out}}$ are low-rank matrices, and $r\ll\min(d,k)$ is the bottleneck rank.
 The product $AB\in\mathbb{R}^{d\times k}$ serves as a low-rank approximation of the quantization error.
-ZeroQuant-v2 [14] and LQER [18] reconstruct the weight quantization error through SVD-based low rank approximation.
-LQER computes the low rank matrices $A,B$ by using SVD composition to the quantization error matrix $\Delta W$:
-$\Delta W=U\Sigma V^{T}\approx\underbrace{U_{k}}_{A_{k}}\cdot\underbrace{\Sigma_{k}V_{k}^{T}}_{B_{k}}=A_{k}B_{k}$ (9)
-where $U\in\mathbb{R}^{m\times n}$, $V\in\mathbb{R}^{n\times n}$, $\Sigma\in\mathbb{R}^{m\times n}$ and $U_{k}\in\mathbb{R}^{m\times k}$, $V_{k}\in\mathbb{R}^{n\times k}$, $\Sigma_{k}\in\mathbb{R}^{k\times k}$.
-Based on LQER, L2QER, which is proposed in LQER, scales the $\Delta W$ before applying SVD, and inversely apply scales in $A_{k}$.
-This operation aims to more precisely approximated the quantization error corresponding to the salient weights.
-The process can be formulated as:
-$\Delta\hat{W}=S\Delta W=U\Sigma V^{T}\approx\underbrace{U_{k}}_{A_{k}}\cdot\underbrace{\Sigma_{k}V_{k}^{T}}_{B_{k}}=A_{k}B_{k}$ (10)
-$\hat{A_{k}}=S^{-1}A_{k},\hat{B_{k}}=Bk$ (11)
-In addition to the aforementioned approaches, CALDERA [16] and QERA [15] provide analyses from different perspectives, deriving the optimal low-rank terms for quantization error reconstruction.
-In contrast to approaches that introduce low-rank branches externally, QwT [21] integrates lightweight linear layers within each decoder layer block of LLMs, directly compensate quantization errors, offering new insights for quantization error mitigation.
-2.3 Symmetric vs. asymmetric
+Table 7 illustrates the computation process of the low-rank terms in some quantization methods.
+Methods Low rank terms $A,B$ computation
+ZeroQuant-v2 [14] $\text{SVD}(dq(W_{q})-W)$
+LQER [16] $\text{SVD}(dq(W_{q})-W)$
+L2QER [16] $\text{SVD}(S(dq(W_{q})-W)),A:=S^{-1}A$
+CALDERA [21] LPLR factorization [21]
+QERA [22] $\text{SVD}(R^{\frac{1}{2}}_{X\!X}(dq(W_{q})-W)),A:=(R^{\frac{1}{2}}_{X\!X})^{-1}A$
+Table 7: Comparison of low-rank decomposition compensation quantization methods. $A,B$ represents the low rank terms. $R_{X\!X}$ denotes the autocorrelation matrix of the input $X$.
+ZeroQuant-v2 [14] and LQER [16] apply SVD composition to reconstruct the quantization error matrix $E=W-W_{q}$.
+Based on LQER, L2QER, which is proposed in LQER, scales the $E$ before applying SVD heuristically, and inversely applies scales in $A_{k}$.
+This operation aims to more precisely approximate the quantization error corresponding to the salient weights.
+In addition to the aforementioned approaches, CALDERA [21] and QERA [22] derive the optimal low-rank terms for quantization error mitigation by framing it as an optimization problem that minimizes the squared error between each layer’s output in Eq.
+(5) and its original output.
+In contrast to approaches that introduce low-rank branches externally, QwT [23] integrates lightweight linear layers within each decoder layer block of LLMs, directly compensating quantization errors, offering new insights for quantization error mitigation.
+2.2 Settings of quantization
+2.2.1 Symmetric vs. asymmetric
 Uniform quantizers can be broadly categorized into symmetric and asymmetric schemes, depending on how the zero point is handled during the mapping from floating-point to integer values.
@@
 Symmetric quantization forces the zero point to align with the floating point, resulting in a quantization range that is symmetric with respect to zero.
-The quantization process is typically defined as:
-$X_{q}=\text{round}\left(\frac{X}{\Delta}\right),\quad\Delta=\frac{\max(|X|)}{2^{N}-1}$ (12)
-where the floating-point tensor $X$ is scaled by a parameter $\Delta$, which is computed from the distribution of $X$.
-Eq.
-(12) presents a commonly used Min-Max method for computing the quantization step size $\Delta$, where $N$ denotes the bit-width of the quantization precision.
-The dequantization process can be formulated as:
-$\text{dequant}(X)=X_{q}\Delta$ (13)
-$\text{dequant}(Y_{q})=X_{q}W_{q}\Delta_{X}\Delta_{W}=Y_{q}\Delta_{X}\Delta_{W}$ (14)
-It is evident that symmetric quantization enables a very simple dequantization process.
-For linear layers, dequantization can be efficiently performed by multiplying the quantized output with the product of the scale factors for the weights and activations like Eq.
-(14).
-This simplifies computation and is often preferred in hardware implementations.
-However, symmetric quantization may lead to suboptimal representation when the data distribution is not centered around zero.
 Asymmetric quantization allows the zero point to shift dynamically according to the minimum and maximum values of the data.
 This enables a tighter fit to the data distribution, especially when it is skewed, at the cost of slightly more complex computation.
+Symmetric quantization enables a very simple dequantization process.
+For linear layers, dequantization can be efficiently performed by multiplying the quantized output by the product of the scale factors for the weights and activations This simplifies computation and is often preferred in hardware implementations.
+However, symmetric quantization may lead to suboptimal representation when the data distribution is not centered around zero.
 Asymmetric quantization incurs additional computational cost during dequantization, while it often outperforms symmetric quantization in terms of accuracy.
-The dequantization process can be formulated as:
-$\text{dequant}(X)=(X_{q}-Z)\Delta$ (15)
-$\text{dequant}(Y_{q})=(X_{q}-Z_{X})\Delta_{X}\cdot(W_{q}-Z_{W})\Delta_{W}$
-$=Y_{q}\Delta_{X}\Delta_{W}+(Z_{X}W_{q}+Z_{W}X_{q}+Z_{X}Z_{W})\Delta_{X}\Delta_{W}$ (16)
 As such, the selection of quantization symmetry should be made with consideration of the specific use case and performance-complexity trade-offs.
-2.4 Granularity of quantization
-The granularity of quantization often largely effects the low-bit quantization performance.
-Commonly adopted granularities include per-tensor, per-channel or per-token, and per-group, as illustrated in Fig.
+2.2.2 Granularity of quantization
+The granularity of quantization often largely affects the low-bit quantization performance.
+Commonly adopted granularities include per-tensor, per-channel, or per-token, and per-group, as illustrated in Fig.
 4.
 Figure 4: Granularity of quantization: (1) Per-tensor quantization; (2) Per-channel quantization; (3) Per-group quantization.
-In per-tensor quantization, a single set of affine transformation parameters is applied to the entire tensor, with the scale ($\Delta$) and zero point ($Z$) computed from the tensor minimum and maximum values.
+In per-tensor quantization, a single set of affine transformation parameters is applied to the entire tensor, with the scale ($\Delta$) and zero point ($Z$) computed from the tensor’s minimum and maximum values.
 A finer-grained alternative is per-channel quantization for weights and per-token quantization for activations, where each input channel of the weights or each token of the activations shares the same quantization parameters.
@@
 Finer granularity typically leads to improved performance but also incurs higher memory overhead due to the need to store more quantization parameters.
-Besides, as QServe [17] mentioned, per-group quantization will also raise computational cost in hardware level.
+Besides, as QServe [15] mentioned, per-group quantization will also raise computational cost at the hardware level.
 Therefore, while finer granularity can enhance accuracy, it may also reduce the overall efficiency of the system.
@@
 Inspired by this, subsequent papers like FlatQuant [12], ResQ [17], and OSTQuant [4] have also adopted per-channel granularity for quantization.
-2.5 MXFP4 data format
-Recently, the NVIDIA’s GeForce RTX 50 Series based on NVIDIA’s Blackwell architecture [22] is released, which has enabled very low-precision operations using FP4 variants like MXFP4 [23, 24] and NVFP4 [25].
-MXFP4 stores each value in an E2M1 format, consisting of one sign bit, one mantissa bit, and two exponent bits.
-Unlike the INT4 format, MXFP4 employs a fixed granularity, utilizing symmetric per-group quantization with a group size of 32.
-Each group is associated with an E8M0 scaling factor, and this rescaling process is hardware-integrated.
-The value of an encoding MXFP4 data is represented as:
+2.3 FP4 Quantization and Variant Formats
+Recently, NVIDIA’s GeForce RTX 50 Series based on NVIDIA’s Blackwell architecture [24] has been released, which has enabled very low-precision operations using FP4 variants like MXFP4 [25, 26] and NVFP4 [27].
+Both MXFP4 and NVFP4 store each value in an E2M1 format, which consists of one sign bit, one mantissa bit, and two exponent bits.
+Unlike the INT4 format, they adopt a fixed granularity and use symmetric per-group quantization, with MXFP4 employing a group size of 32 and NVFP4 using a finer granularity of 16.
+Each group is associated with a scaling factor, and this rescaling process is hardware-accelerated.
+MXFP4 uses an FP8 (E8M0) format, while NVFP4 uses an FP8 (E4M3) format.
+Additionally, NVFP4 incorporates an extra FP32 scaling factor per tensor to extend its representable range.
+The value of an encoding FP4 (E2M1) data is represented as:
 $\text{value}=\begin{cases}(-1)^{S}\times 2^{1-\text{bias}}\times(0+2^{-m}\times M)&\text{if }E=0\\ (-1)^{S}\times 2^{E-\text{bias}}\times(1+2^{-m}\times M)&\text{if }E>0\end{cases}$ (6)
-where $S,E,M$ corresponds to the sign,exponent and mantissa fields, respectively.
-The bias is the exponent bias which is set to 1 for MXFP4.
-So the MXFP4 encoding values are [-6, -4, -3, -2, -1.5, -1, -0.5, 0, 0.5, 1, 1.5, 2, 3, 4, 6].
-The quantization process of MXFP4 is similar to the integer quantization process.
-First, we compute the scaling factor, using the same way as symmetric quantization situation mentioned above that computes from the maximum value of a group of values.
-Then the scaling factors need to be cast to E8M0 format.
-After scaling, the values are quantized to the nearest MXFP4 encoding values.
-MXFP4 is enable to better handle long-tail distributions compared to INT4, as its values are non-uniformly distributed.
+where $S,E,M$ correspond to the sign, exponent, and mantissa fields, respectively.
+The bias is the exponent bias, which is set to 1.
+So the FP4 (E2M1) encoding values are [-6, -4, -3, -2, -1.5, -1, -0.5, 0, 0.5, 1, 1.5, 2, 3, 4, 6].
+FP4 can better handle long-tail distributions compared to INT4, as its values are non-uniformly distributed.
 In LLMs, both weights and activations approximately follow normal distributions, with activations particularly exhibiting long-tail characteristics.
@@
 A clear observation is that under INT4 quantization, values with larger absolute magnitudes represent fewer data points, whereas MXFP4 more fully utilizes the representational capacity of 4-bit precision.
-Therefore, MXFP4 is more suitable than INT4 as a quantization data type.
+Therefore, FP4 is more suitable than INT4 as a quantization data type.
 (a) Full-precision weight
@@
 3 Extensive Evaluations
-Quantization settings We conduct experiments using W4A4 precision, comparing the results with the origin precision (BF16 or FP16).
+Quantization settings We conduct experiments using W4A4 precision, comparing the results with the original precision (BF16 or FP16).
 In the main results, we use per-token or per-group (group size=128) asymmetric quantization for activations.
@@
 To calibrate the GPTQ [8] quantization and the low-rank compensation branch, we use 128 samples from WikiText-2 with a sequence length of 2048.
-The training of the learnable orthogonal and scaling matrices follows OSTQuant [4] procedure, using 1,000 samples from WikiText-2 and optimized with RiemannAdam.
-For the low-rank branch, we set the rank k=32 and store the low-rank matrices in FP16.
+The training of the learnable orthogonal and scaling matrices follows the OSTQuant [4] procedure, using 1,000 samples from WikiText-2 and optimized with RiemannAdam.
+For the low-rank decomposition, we set the rank k=32 and store the low-rank matrices in FP16.
 For the non-trainable scaling matrix, we follow the SmoothQuant [6] calibration scheme with $\alpha$=0.5.
@@
 Evaluate Methods Channel Group-128
-DN EC Wiki2‑PPL($\downarrow$) 0‑shot9 Avg.($\uparrow$) Wiki2‑PPL($\downarrow$) 0‑shot9 Avg.($\uparrow$)
+PT QEM Wiki2‑PPL($\downarrow$) 0‑shot9 Avg.($\uparrow$) Wiki2‑PPL($\downarrow$) 0‑shot9 Avg.($\uparrow$)
 BF16 9.76 54.78 9.76 54.78
-rotation - 17.83 45.24 14.84 47.51
+Rotation RTN 17.83 45.24 14.84 47.51
 Rotation Low-rank 17.16 45.47 14.5 47.04
@@
 Rotation GPTQ+Low-rank 12.81 48.35 11.61 50.49
-rotation+scaling - 14.88 46.41 13.3 49.22
+Rotation+Scaling RTN 14.88 46.41 13.3 49.22
 Rotation+Scaling Low-rank 14.58 46.76 13.07 49.95
@@
 Table 8: Evaluation of Quantization Method Combinations, in which rotation or scaling is optimized, in terms of Perplexity on WikiText2 and averaged accuracy on nine Zero-Shot tasks across per-channel and per-group settings of LlaMA-3.2-1B.
+PT and QEM represent pre-quantization transformation and quantization error mitigation, respectively.
 Evaluate Methods Channel Group-128
-DN EC Wiki2‑PPL($\downarrow$) 0‑shot9 Avg.($\uparrow$) Wiki2‑PPL($\downarrow$) 0‑shot9 Avg.($\uparrow$)
+PT QEM Wiki2‑PPL($\downarrow$) 0‑shot9 Avg.($\uparrow$) Wiki2‑PPL($\downarrow$) 0‑shot9 Avg.($\uparrow$)
 BF16 9.76 54.78 9.76 54.78
-rotation 20.31 43.26 15.18 46.62
+Rotation RTN 20.31 43.26 15.18 46.62
 Rotation Low‑rank 19.84 43.09 14.83 46.87
@@
 Rotation GPTQ+Low‑rank 13.28 48.37 11.98 50.48
-rotation+scaling 20.96 42.87 16.25 45.01
+Rotation+Scaling RTN 20.96 42.87 16.25 45.01
 Rotation+Scaling Low‑rank 21.97 43.07 15.55 45.46
@@
 Table 9: Evaluation of Quantization Method Combinations, in which rotation or scaling is non-optimized, in terms of Perplexity on WikiText2 and averaged accuracy on nine Zero-Shot tasks across per-channel and per-group settings of LlaMA-3.2-1B.
-We select two of the most widely-used pre-quantization transformation techniques, scaling and rotation, and combined them with two quantization error mitigation methods: GPTQ and low-rank branch compensation.
-A series of experiments are conducted to evaluate these combinations.
+PT and QEM represent pre-quantization transformation and quantization error mitigation, respectively.
+We select two of the most widely used pre-quantization transformation techniques, scaling and rotation, and combine them with two quantization error mitigation methods: GPTQ and low-rank decomposition compensation.
+A series of experiments is conducted to evaluate these combinations.
 Table 8 presents the results of using optimized scaling and rotation, illustrating the performance of different method combinations.
 Table 9 presents the results obtained with randomly initialized rotations and calibrated scaling factors.
-Our experiments reveal that the best performance is achieved when both optimized scaling and rotation are applied in conjunction with GPTQ and the low-rank branch.
-This configuration yields a perplexity of 11.73 in per-channel quantization for Llama-3.2-1B model, which is close to the BF16 baseline of 9.76.
+Our experiments reveal that the best performance is achieved when both optimized scaling and rotation are applied in conjunction with GPTQ and the low-rank decomposition.
+This configuration yields a perplexity of 11.73 in per-channel quantization for the Llama-3.2-1B model, which is close to the BF16 baseline of 9.76.
 Between the two quantization error mitigation methods, GPTQ clearly plays a more significant role than low-rank adaptation.
@@
 In addition, using both optimized rotation and scaling improves performance over using only optimized rotation.
-However, when employing non-optimized rotation and scaling, rotation alone tends to yield better results than the combination, which indicates that calibrated scaling may not performs well together with rotation in W4A4 settings.
-Model precision Wiki2 PPL($\downarrow$) Zero-Shot Accuracy($\uparrow$)
+However, when employing non-optimized rotation and scaling, rotation alone tends to yield better results than the combination, which indicates that calibrated scaling may not perform well together with rotation in W4A4 settings.
+Model precision
+Wiki2
+PPL($\downarrow$)
+Zero-Shot Accuracy($\uparrow$)
 General Sentiment Math Reasoning QA Avg.
 MMLU SST2 MathQA PiQA HellaS WinoG BoolQ Lambada Arc-C
-Llama- 3.2-1B fp16 9.76 36.76 66.86 28.94 74.65 63.67 60.38 63.67 62.1 36.01 54.78
+Llama-
+3.2-1B
+fp16 9.76 36.76 66.86 28.94 74.65 63.67 60.38 63.67 62.1 36.01 54.78
 w4a4 11.73 29.21 55.28 25.43 69.15 57.93 56.43 59.02 56.78 32.34 49.06
-Llama- 3.2-3B fp16 7.82 54.06 73.62 34.47 77.48 73.58 69.46 72.97 70.08 46.08 63.53
+Llama-
+3.2-3B
+fp16 7.82 54.06 73.62 34.47 77.48 73.58 69.46 72.97 70.08 46.08 63.53
 w4a4 8.9 46.7 68.58 30.55 73.99 69.54 65.43 68.75 65.32 41.55 58.93
-Qwen- 2.5-3B fp16 8.03 65.13 90.37 37.15 78.73 73.55 68.75 77.58 66.8 47.35 67.27
+Qwen-
+2.5-3B
+fp16 8.03 65.13 90.37 37.15 78.73 73.55 68.75 77.58 66.8 47.35 67.27
 w4a4 9.22 56.32 69.15 32.5 74.16 67.52 62.12 65.41 57.77 42.83 58.64
-Table 5: Comparision of s ymmetric and asymmetric quantization in Terms of Perplexity on WikiText2 and averaged accuracy on Zero-Shot tasks of LlaMA-3.2-1B
-In Table 5, we further conduct experiments on different models: LLaMA-3.2-1B, LLaMA-3.2-3B, and Qwen-2.5-3B, using the optimized scaling and rotation, and combined them with GPTQ and an additional low-rank branch.
+Table 10: Comparison of symmetric and asymmetric quantization in Terms of Perplexity on WikiText2 and averaged accuracy on Zero-Shot tasks of LlaMA-3.2-1B
+In Table 10, we further conduct experiments on different models: LLaMA-3.2-1B, LLaMA-3.2-3B, and Qwen-2.5-3B, using the optimized scaling and rotation, and combined them with GPTQ and an additional low-rank decomposition.
 3.2 Symmetric vs. asymmetric
@@
 Arc-Challenge Score(%) 31 31.7 33.3 33.7 31.9 29.9 32.3 32.7
-Table 6: Comparision of symmetric and asymmetric quantization in terms of Perplexity on WikiText2 and Arc [32] accuracy of LlaMA-3.2-1B
+Table 11: Comparison of symmetric and asymmetric quantization in terms of Perplexity on WikiText2 and Arc [34] accuracy of LlaMA-3.2-1B
 In Table 11, we evaluate the performance of applying symmetric or asymmetric quantization to weights and activations, respectively.
@@
 Therefore, an effective configuration is to use symmetric quantization for weights and asymmetric quantization for activations.
-This setup not only achieves much better performance than fully symmetric quantization, but also reduces the computational overhead associated with dequantization process, compared to fully asymmetric quantization for both weights and activations.
+This setup not only achieves much better performance than fully symmetric quantization, but also reduces the computational overhead associated with the dequantization process, compared to fully asymmetric quantization for both weights and activations.
 3.3 Granularity
@@
 The experimental results reveal that finer quantization granularity yields superior model performance (lower perplexity), albeit at the expense of elevated storage requirements for quantization parameters.
-This trade-off stems primarily from the FP16-format scaling factors, whose storage overhead scales inversely with group size.
+This trade-off stems primarily from the scaling factors, whose storage overhead scales inversely with group size.
 Quantitative analysis of these additional bit costs is presented in Table 12, which compares the average bit overhead introduced by scaling factors under symmetric weight quantization.
-Figure 6: Comparision of different group sizes of quantization in terms of Perplexity on WikiText2 and averaged accuracy on Zero-Shot tasks of LlaMA-3.2-1B
+Figure 6: Comparison of different group sizes of quantization in terms of Perplexity on WikiText2 and averaged accuracy on Zero-Shot tasks of LlaMA-3.2-1B
 Granularity group-32 group-64 group-128 group-256 group-512
-Average Extra bits 0.5 0.25 0.125 0.0625 0.03125
-Table 7: Average extra bits introduced by scaling factors of different granularities
-3.4 MXFP4 vs. INT4
-Extending experiments to the MXFP4 format, we follow MXFP4 configurations implementing symmetric quantization with group size 32.
-For activations, we use a clipping threshold of 3/4.
-For weights, we follow the experiments of INT4 quantization, using a linear search over the squared error to find a best clipping ratio.
-First, we apply no extra pre-quantization transformation or quantization error mitigation process, conducting experiments using only RTN (round to nearest).
-Table 8 demonstrates that, MXFP4 shows much better performance than INT4 (per-channel).
-Error compensation Data Format
-INT4 (channel) INT4 (g-32) MXFP4
-- 923.72 14.19 15.19
-GPTQ 1007.33 14.73 13.18
-low-rank 723.54 14.04 15.59
-GPTQ+low-rank 578.61 12.48 12.72
-Table 8: Evaluation of MXFP4 format quantization without using pre-quantization transformation, in terms of Perplexity on WikiText2 of LlaMA-3.2-1B.
-Evaluate Methods WikiText2PPL($\downarrow$)
-Data Normalization Error Compensation INT4(channel) MXFP4
-scaling GPTQ 134.85 13.06
-scaling low-rank 165.4 14.83
-scaling GPTQ+low-rank 124.56 12.64
-optimized rotation+scaling GPTQ 11.8 12.61
-optimized rotation+scaling low-rank 14.58 15.53
-optimized rotation+scaling GPTQ+low-rank 11.73 12.29
-Table 9: Evaluation of MXFP4 format quantization, in terms of Perplexity on WikiText2 of LlaMA-3.2-1B.
-Further, we conduct experiments using both pre-quantization transformation and quantization error mitigation techniques.
-Results in Table 9 demonstrate that both pre-quantization transformation and quantization error mitigation techniques achieve competitive performance under MXFP4 quantization, confirming the direct applicability of INT4 quantization methods to FP4.
-Crucially, scaling-based approaches proved more effective than rotation-based methods for MXFP4 –- a reversal of their relative efficacy observed in INT4 quantization scenarios.
+Average Extra bits
+1 0.5 0.25 0.125 0.0625
+Table 12: Average extra bits introduced by FP32 scaling factors of different granularities
+3.4 FP4 vs. INT4
+Extending our experiments to the FP4 format, we compare several variants of FP4, including MXFP4, NVFP4, and naive FP4 implementation with varying granularities and scaling factor formats.
+For per-channel quantization, we employ asymmetric quantization for activations and symmetric quantization for weights.
+To ensure a fair comparison with MXFP4 and NVFP4, all other experiments using per-group quantization adopt symmetric quantization for both activations and weights.
+First, we apply no extra pre-quantization transformation, conducting experiments using only quantization error mitigation methods.
+Table 13 demonstrates that FP4 (per-channel) shows much better performance than INT4 (per-channel).
+We believe this is because, in the absence of the pre-quantization transformation, FP4 better aligns with the data distribution, thus achieving superior results.
+For finer granularity scenarios, FP4 and INT4 show very small differences, and both the format of the scaling factor and the size of the group significantly impact performance.
+Compared to MXFP4, NVFP4 achieves better performance due to its smaller group size and more precise scaling factor format.
+With the E4M3 variant of FP8, NVFP4 can encode scaling factors with fractional precision.
+By contrast, MXFP4 uses the E8M0 format, which restricts scale factors to powers of two, resulting in much coarser quantization.
+Moreover, NVFP4 employs a two-level micro-block scaling strategy: it first applies a per-tensor scaling factor (FP32) to compensate for the smaller range that E4M3 scales can represent.
+NVFP4 performs better than MXFP4, showing a lower risk of noticeable accuracy drop, matching NVIDIA’s official conclusions [40].
+Quantization
+error
+mitigation
+Data Format
+INT4
+channel, FP32
+FP4
+channel, FP32
+INT4
+g16, FP32
+FP4
+g16, FP32
+MXFP4
+g32, E8M0
+FP4
+g32, E4M3
+FP4
+g16, E8M0
+NVFP4
+g16, E4M3
+RTN 256.97 135.68 11.75 11.34 15.17 12.04 14.52 11.54
+GPTQ 161.04 101.36 11.47 10.87 13.24 11.36 12.77 10.98
+low-rank 179.87 132.14 11.69 11.32 15.56 11.98 14.67 11.48
+GPTQ+low-rank 139.31 96.33 10.92 10.76 12.69 11.18 12.25 10.83
+Table 13: Evaluation of INT4 and FP4 format quantization without using pre-quantization transformation, in terms of Perplexity on WikiText2 of LlaMA-3.2-1B.
+Channel means quantization granularity is per-channel; g16 and g32 denote per-group quantization with group sizes of 16 and 32, respectively.
+FP32, E8M0, and E4M3 refer to the format of the scaling factor.
+And for E4M3, the scaling strategy follows NVFP4’s approach, employing a two-level scaling scheme.
+Quantization
+error
+mitigation
+Data Format
+NVFP4
+W:E4M3, A:E4M3
+FP4
+W:E2M1, A:E4M3
+FP4
+W:E2M1, A:E2M1
+FP4
+W:INT4, A:E4M3
+FP4
+W:INT4, A:INT4
+RTN 11.54 14.10 4130 14.66 4777
+GPTQ 10.98 13.38 3618 17.96 3803
+Low-rank 11.48 13.75 4057 14.43 5356
+GPTQ+Low-rank 10.83 11.38 3650 11.59 4124
+Table 14: Evaluation of FP4 format quantization using low bit scaling factor and group size 16, in terms of Perplexity on WikiText2 of LlaMA-3.2-1B.
+E4M3, E2M1, and INT4 refer to the format of the scaling factor.
+W and A denote weight and activation, respectively.
+All of these follow NVFP4’s approach, employing a two-level scaling scheme.
+We further reduce the bit-width of the scaling factor to 4 bits to evaluate its effectiveness under such a low-bit setting.
+As shown in Table 14, applying 4-bit scaling factors to both weights and activations leads to very poor performance.
+When applied only to weights, the performance remains moderate and is worse compared with NVFP4.
+Evaluate Methods Wiki2‑PPL($\downarrow$)
+Pre-quantization
+transformation
+Quantization error
+mitigation
+INT4
+channel, FP32
+MXFP4
+g32, E8M0
+NVFP4
+g16, E4M3
+scaling GPTQ 134.85 13.24 10.97
+scaling low-rank 165.4 14.91 11.36
+scaling GPTQ+low-rank 124.56 12.58 10.81
+optimized rotation+scaling GPTQ 11.8 12.37 11.04
+optimized rotation+scaling low-rank 14.58 15.43 12.57
+optimized rotation+scaling GPTQ+low-rank 11.73 12.29 11.01
+Table 15: Evaluation of INT4, MXFP4, and NVFP4 using both pre-quantization transformation and quantization error mitigation techniques, in terms of Perplexity on WikiText2 of LlaMA-3.2-1B.
+Channel means quantization granularity is per-channel; g16 and g32 denote per-group quantization with group sizes of 16 and 32, respectively.
+FP32, E8M0, and E4M3 refer to the format of the scaling factor.
+Notably, rotation-based methods yield smaller improvements for MXFP4 and NVFP4 than for INT4 (per-channel).
+As argued in AMXFP4 [41], grouping activation tensors into small micro-scaled units mitigates outliers like rotation.
+For MXFP4 and NVFP4 formats, the group size is extremely small, so data rotation contributes little and may even lead to degradation.
 4 Conclusions
 In this paper, we have presented an experimental evaluation of Large language model quantization.
-We systematically decouple the commonly used quantization methods into two key steps: pre-quantization transformation and quantization error mitigation.
-After a comprehensive review of existing approaches, we choose scaling and rotation in pre-quantization transformation step, GPTQ and low-rank compensation in quantization error mitigation step to further conduct experimental evaluations on the same ground.
-We found that uses the optimized rotation and scaling, together with two compensation methods demonstrating the best performance.
+We systematically decouple the commonly used quantization into two key steps: pre-quantization transformation and quantization error mitigation.
+After a comprehensive review of existing approaches, we choose scaling and rotation in pre-quantization transformation step, GPTQ, and low-rank compensation in the quantization error mitigation step to further conduct experimental evaluations on the same ground.
+We found that using the optimized rotation and scaling, together with two compensation methods, demonstrates the best performance.
 For quantization settings, asymmetric quantization and smaller group sizes yield better results.
-For MXFP4 quantization, we can adopt methods used in INT4 quantization.
-However, using scaling-based pre-quantization transformation demonstrates superior performance compared to INT4.
-To achieve optimal performance under MXFP4 quantization, further investigation is required to explore its full potential and limitations.
+And using asymmetric quantization on activations shows more benefits on performance compared with weights.
+For FP4 quantization, we can adopt methods used in INT4 quantization.
+However, rotation-based pre-quantization transformation yields no observable more performance gains for MXFP4 and NVFP4 than scaling-based methods.
+And we find that the format of scaling factors demonstrates a large effect.
+For FP8 format scaling factor, FP8 (E4M3) with an additional per-tensor FP32 scaling factor is better than FP8 (E8M0).
+For small group sizes, using a 4-bit scaling factor for weights is also feasible, albeit with some loss in accuracy.
+To achieve optimal performance under FP4 quantization, further investigation is required to explore FP4’s full potential and limitations.

### v3

@@
 Though many papers have reported breakthrough performance, they may not conduct experiments on the same ground since one quantization method solution usually contains multiple components.
-In addition, it is also important to analyze the connections among numerous existing methods to gain in-depth understanding of this field.
-To bridge these gaps, in this paper, we timely conduct an extensive review of the state-of-the-art methods and also perform comprehensive evaluations on the same ground to ensure fair comparisons of these methods.
+In addition, it is also important to analyze the connections among numerous existing methods to gain an in-depth understanding of this field.
+To bridge these gaps, in this paper, we conduct an up-to-date and extensive review of the state-of-the-art methods and also perform comprehensive evaluations on the same ground to ensure fair comparisons of these methods.
 To our knowledge, this fair and extensive investigation remains critically important yet significantly underexplored in the research community.
 To better understand the connections, first, we decouple the published quantization methods into two steps: pre-quantization transformation and quantization error mitigation.
-We define the former as a preprocessing step applied before quantization to reduce the impact of outliers, making the data distribution flatter and more suitable for quantization.
-Quantization error mitigation, on the other hand, involves techniques that offset the errors introduced during quantization, thereby enhancing model performance.
+We define the former as a preprocessing step applied before quantization to reduce the impact of outliers, making the data distribution more suitable for quantization.
+Quantization error mitigation involves techniques that offset the errors introduced during quantization, thereby enhancing model performance.
 Second, we extensively evaluate and analyze the impact of different settings of quantization methods, including granularity and symmetry.
-Third, we also timely analyze and evaluate the latest MXFP4 and NVFP4 data format and its performance.
+Third, we also analyze and evaluate the latest MXFP4 and NVFP4 data formats and their performance.
 Our experimental results first demonstrate that the optimized rotation and scaling yield the best performance for pre-quantization transformation, and combining low-rank compensation with GPTQ occasionally outperforms using GPTQ alone for quantization error mitigation.
@@
 We also verified that applying asymmetric quantization to weights yields much smaller benefits compared to applying it to activations, which is consistent with the common practice of using asymmetric quantization primarily to activations.
-Third, we explore the potential of the latest FP4 quantization and find that the format and precision of scaling factors have a significant impact on the performance.
+Third, we explore the potential of the latest FP4 quantization and find that the format and precision of scaling factors have a significant impact on performance.
 We also show the performance of the optimal pre-quantization transformation strategy for INT4 on FP4, revealing that rotation-based methods demonstrate little improvement in performance for MXFP4 and NVFP4, inspiring further investigation in this field.
-keywords:
-Model Quantization, Post-training Quantization , LLM
-1 Introduction
+I Introduction
 Large language models (LLMs) have achieved remarkable performance on diverse tasks and are increasingly being applied in various scenarios.
-Nonetheless, LLMs inference process incurs significant memory and computational overhead, leading to high inference latency and posing challenges for computations on source-limited GPUs.
-Post-training quantization (PTQ) is a promising solution to reduce the memory and computational cost of LLMs, enhancing their inference efficiency.
-LLMs quantization has been studied in a wide spectrum of precision configurations.
-Representative settings include W4A4 [1, 2, 3, 4], W8A8 [5, 6, 7] and W4A16 [8, 9, 10, 11]: W4A4 quantizes both weights and activations to 4-bit precision (INT4); W8A8 and W4A16 quantize weights to INT8 and INT4 and activations to INT8 and FP16/BF16 respectively.
+Nonetheless, LLMs’ inference process incurs significant memory and computational overhead, leading to high inference latency and posing challenges for computations on resource-limited GPUs.
+Quantization is a promising solution to reduce the memory and computational cost of LLMs, enhancing their inference efficiency.
+In recent years, the quantization techniques have developed rapidly.
+Various approaches have been proposed, including Quantization-Aware Training (QAT), Post-Training Quantization (PTQ), and hybrid methods that combine quantization with knowledge distillation or pruning.
+In this work, we focus solely on quantization, without considering hybrid approaches.
+In addition, we focus on PTQ, because it is more widely used than QAT in real applications.
+Additionally, LLMs’ quantization has been studied in a wide spectrum of precision configurations.
+Representative settings include W4A4 [43, 32, 22, 13], W8A8 [8, 38, 37] and W4A16 [10, 6, 35, 19] and mixed-precision settings such as W4A8 [20]: W4A4 quantizes both weights and activations to 4-bit precision (INT4); W8A8 and W4A16 quantize weights to INT8 and INT4 and activations to INT8 and FP16/BF16 respectively; Mixed-precision methods aim to improve accuracy by using higher bit-widths for parts that are more difficult to quantize [43].
 Among these three settings, W4A4 is currently the main focus of research, as it holds great potential for improving memory efficiency and inference speed.
 However, applying lossless W4A4 quantization to LLMs remains a challenging problem.
-Thus, we concentrate our efforts on exploring W4A4 quantization.
-In this paper, we mainly investigate three dimensions.
-First, we find that existing methods are often combinations of multiple techniques, since using a single technique alone typically fails to achieve excellent quantization performance.
+To make our evaluation more cutting-edge and provide the research community with up-to-date insights, we select W4A4 as the primary precision configuration for our study.
+Many papers report breakthrough performance in quantization; however, their findings are often not directly comparable, as quantization often contains multiple configurations.
+It is important to evaluate these methods on the same ground.
+Besides, we find that existing methods are often combinations of multiple techniques, since using a single technique alone typically fails to achieve excellent quantization performance.
 OSTQuant [10] and FlatQuant [34] employ multiple techniques to adjust the weights and activations, including rotation and scaling, and also apply the GPTQ [10], a gradient-based method for mitigating quantization error, to quantize the transformed weights.
 Besides, many techniques share similar transformations, for instance, Quip# [35] and QuaRot [3] both use rotation, also, AWQ [19] and SmoothQuant [38] both use scaling.
-To achieve a better understanding of existing methods and make our evaluation on the same ground, we decompose the quantization process into two key steps: pre-quantization transformation and quantization error mitigation.
+Analysis of the interconnections among diverse methods is critical for a thorough understanding of the field.
+To bridge these gaps, in this paper, we conduct an up-to-date and extensive review of the state-of-the-art methods and also perform comprehensive evaluations on the same ground to ensure fair comparisons of these methods.
+Although several excellent survey papers on quantization exist [17, 14], this work makes more extensive evaluations and categorizes more state-of-the-art quantization methods from different perspectives.
+We mainly investigate three dimensions.
+First, to achieve a better understanding of existing methods and make our evaluation on the same ground, we decompose the quantization process into two key steps: pre-quantization transformation and quantization error mitigation.
 The pre-quantization transformation step refers to preprocessing the data before quantization in order to mitigate the effect of outliers, making the data distribution smoother and more quantization-friendly.
@@
 However, the research on FP4 has not been well explored.
-To bridge this gap, we conduct research on weight-activation FP4 quantization using the popular methods in INT4, showing the great timeliness of our research.
-For FP4 quantization, we also observe that the format and precision of scaling factors have a significant impact on the performance.
-2 Review on Existing Quantization Methods
-In this section, we first review recent quantization algorithms in Section 2.1 in terms of two steps: pre-quantization transformation and quantization error mitigation.
-Table 1 shows our decomposition of the popular quantization methods.
-Section 2.2 focuses on settings, including symmetry and granularity of quantization methods.
-Meanwhile, Table 2 summarizes the settings of existing methods, showing that they are not compared on the same ground.
-Section 2.3 discusses FP4 quantization and its variants, including MXFP4 and NVFP4.
+To bridge this gap, we conduct research on weight-activation FP4 quantization under our two-step quantization framework in INT4, showing the great timeliness of our research.
+For FP4 quantization, we also observe that the format and precision of scaling factors have a significant impact on performance.
+II Review on Existing Quantization Methods
+In this section, we first review recent quantization algorithms in Section II-A in terms of two steps: pre-quantization transformation and quantization error mitigation.
+Table I shows our decomposition of the popular quantization methods.
+Section II-B focuses on settings, including symmetry and granularity of quantization methods.
+Meanwhile, Table II summarizes the settings of existing methods, showing that they are not compared on the same ground.
+Section II-C discusses FP4 quantization and its variants, including MXFP4 and NVFP4.
+TABLE I: Two-step decomposition of quantization methods.
 Methods Pre-quantization Transformation Quantization Error Mitigation
@@
 Outlier Suppression+ [37] Shifting + Scaling RTN
-GPTQ [8] - GPTQ
-ZeroQuant-v2 [14] - Low-rank
+GPTQ [10] GPTQ
+ZeroQuant-v2 [39] Low-rank
 QuIP [6] Rotation GPTQ
@@
 OSTQuant [13] Scaling + Rotation GPTQ
-Table 1: Two steps decomposition of quantization methods.
-2.1 Two-steps of quantization
-2.1.1 Pre-quantization transformation
+TABLE II: Symmetry and granularity of quantization methods. For granularity, T: per-tensor, C: per-channel/token, G: per-group.
+Methods Precision Symmetric Gran.
+W A
+SmoothQuant [38] W8A8 T F T, C
+GPTQ [10] W3A16, W4A16 T G
+ZeroQuant-v2 [39] W2/3/4A16, W4A8 T, F T, F G
+QuIP [6] W2/3/4A16 T G
+AWQ [19] W3A16, W4A16 T G
+Atom [43] W3A3, W4A4 T F G
+OmniQuant [32] W2/3/4A16, W4/6/8A4/6/8 F F C
+Qserve [20] W4A8 T F C
+LQER [41] W4A6, W4A8 T F G
+QuaRot [3] W4A4 T F C, G
+SpinQuant [22] W4A4, W4A8 T F C
+ResQ [31] W4A4 T F C
+FlatQuant [34] W4A4 T F C
+OSTQuant [13] W4A4 T F C
+II-A Two steps of quantization
+II-A1 Pre-quantization transformation
 Pre-quantization transformation is a step before quantization, applying operations such as shifting, scaling, and rotation to process weights and activations.
@@
 1.
-Methods Precision Symmetric Granularity
-W A
-SmoothQuant [6] W8A8 T F T,C
-GPTQ [8] W3A16,W4A16 T - G
-ZeroQuant-v2 [14] W4A16,W3A16,W2A16,W4A8 T,F T,F G
-QuIP [9] W2A16,W3A16,W4A16 T - G
-AWQ [11] W3A16,W4A16 T - G
-Atom [1] W3A3,W4A4 T F G
-OmniQuant [2] W2A16,W3A16,W4A16,W4A4,W6A6,W8A8 F F C
-Qserve [15] W4A8 T F C
-LQER [16] W4A6,W4A8 T F G
-QuaRot [13] W4A4 T F C,G
-SpinQuant [3] W4A4,W4A8 T F C
-ResQ [17] W4A4 T F C
-FlatQuant [12] W4A4 T F C
-OSTQuant [4] W4A4 T F C
-Table 2: Symmetry and granularity of quantization methods. In granularity, T stands for per-tensor, C represents per-channel or per-token, and G means per-group.
-Figure 1: Pre-quantization transformation: a process that transforms activations and weights to make them easier to quantize.
-The left graph illustrates a tensor that is difficult to quantize due to the presence of numerous outliers, while the right graph shows that after the pre-quantization transformation, the tensor becomes more uniformly distributed, making quantization easier.
 Shifting
@@
 To keep mathematical equivalence, an additional term, the product of $T$ and $W\in\mathbb{R}^{C_{in}\times C_{out}}$, is added to the bias $B\in\mathbb{R}^{1\times C_{out}}$.
+Figure 1: Pre-quantization transformation: a process that transforms activations and weights to make them easier to quantize.
+The left graph illustrates a tensor that is difficult to quantize due to the presence of numerous outliers, while the right graph shows that after the pre-quantization transformation, the tensor becomes more uniformly distributed, making quantization easier.
+TABLE III: Comparison of shifting-based quantization methods.
 Methods Shifting Strategy Shifting parameters $T$
@@
 OmniQuant [32] Pre-scaling shifting Gradient-based optimization
-Table 3: Comparison of shifting-based quantization methods.
-As shown in Table 3, Outlier Suppression+ [7] uses calibration to get the vector $T$, while OmniQuant [2] employs learnable channel-wise shifting for outlier suppression, combined with the learnable scaling technique discussed in the next paragraph, and jointly optimizes these two kinds of pre-quantization transformation parameters.
+As shown in Table III, Outlier Suppression+ [37] uses calibration to get the vector $T$, while OmniQuant [32] employs learnable channel-wise shifting for outlier suppression, combined with the learnable scaling technique discussed in the next paragraph, and jointly optimizes these two kinds of pre-quantization transformation parameters.
 Scaling
@@
 where $S$ is a diagonal matrix composed of scaling factors of different channels, which can be obtained in various ways.
-Methods Scaling Strategy Scaling parameters $S$
-SmoothQuant [6] Scaling only $\begin{aligned} &S_{i}=\frac{\max(|X_{i}|)^{\alpha}}{\max(|W_{i}|)^{1-\alpha}}\end{aligned}$
-AWQ [11] Scaling only $\begin{aligned} &S_{i}=|X_{i}|_{avg}^{\alpha}\\ &\alpha^{*}=\arg\min_{\alpha}\mathcal{L}(S_{Xi}^{\alpha})\end{aligned}$
-OS+ [7] Post-shifting scaling $\begin{aligned} &S_{i}=\max\left(1.0,\frac{\max(X_{i}-T_{i})}{t}\right)\\ &t^{*}=\arg\min_{t}\mathcal{L}(S_{i})\end{aligned}$
-OmniQuant [2] Post-shifting scaling Gradient-based optimization
-FlatQuant [12] Pre-rotation scaling Calibration initialization and gradient-based optimization
-OSTQuant [4] Post-rotation scaling Gradient-based optimization
-Table 4: Comparison of scaling-based quantization methods.
-As shown in Table 4, SmoothQuant [6] uses calibration to compute $S$.
+TABLE IV: Comparison of scaling-based quantization methods.
+Methods Strategy Scaling parameters $S$
+SmoothQuant [38] Scaling only $S_{i}=\max(|X_{i}|)^{\alpha}/\max(|W_{i}|)^{1-\alpha}$
+AWQ [19] Scaling only $S_{i}=|X_{i}|_{\text{avg}}^{\alpha}$ $\alpha^{*}=\arg\min_{\alpha}\mathcal{L}(S_{Xi}^{\alpha})$
+OS+ [37] Post-shift $S_{i}=\max(1.0,\frac{\max(X_{i}-T_{i})}{t})$ $t^{*}=\arg\min_{t}\mathcal{L}(S_{i})$
+OmniQuant [32] Post-shift Gradient-based optimization
+FlatQuant [34] Pre-rot. Calibration init. & Grad-based
+OSTQuant [13] Post-rot. Gradient-based optimization
+As shown in Table IV, SmoothQuant [38] uses calibration to compute $S$.
 AWQ [19], as a weight-only quantization, approaches the problem from the perspective of protecting salient weights.
@@
 Outlier Suppression+ [37] also uses grid searching optimization to find the optimal scaling after applying shifting.
-Unlike adopting the hand-crafting parameters in calibration, Outlier Suppression+ [7] uses grid searching optimization to find the optimal scaling after applying shifting.
 OmniQuant [32] optimizes the quantization parameters via gradient-based methods to minimize the block-wise quantization error.
@@
 To preserve the model’s output, the weight $W$ is simultaneously transformed by $O^{\top}$, which is the transpose of $O$.
-Methods Rotation Strategy Rotation parameters $O$
+TABLE V: Comparison of rotation-based quantization methods.
+Methods Strategy Rotation parameters $O$
 QuIP [6] Rotation only Random orthogonal matrices
@@
 SpinQuant [22] Rotation only Gradient-based optimization
-FlatQuant [12] Post-scaling Rotation Gradient-based optimization
-OSTQuant [4] Pre-scaling Rotation Gradient-based optimization
-Table 5: Comparison of rotation-based quantization methods.
-As shown in Table 5, QuIP [9] first introduced rotation with random orthogonal matrices to reduce incoherence in weight-only quantization.
+FlatQuant [34] Post-scaling Gradient-based optimization
+OSTQuant [13] Pre-scaling Gradient-based optimization
+As shown in Table V, QuIP [6] first introduced rotation with random orthogonal matrices to reduce incoherence in weight-only quantization.
 QuIP# [35] replaces this with random Hadamard matrices for greater efficiency.
@@
 Rotation is a highly effective pre-quantization transformation technique for reducing outliers in low-bit quantization, and it has also been widely adopted in many other quantization methods together with other pre-quantization transformation techniques to further improve performance (e.g., ResQ [31], QServe [20], FlatQuant [34], and OSTQuant [13], etc).
-2.1.2 Quantization error mitigation
+II-A2 Quantization error mitigation
 The quantization error mitigation step employs techniques that mitigate the error introduced by quantization by compensating for it.
@@
 Existing methods that employ self-compensation are generally based on analyzing the loss error introduced by quantization and make use of its Hessian-based quadratic approximation, which can be formulated as:
-$\Delta\mathcal{L}=\mathcal{L}(dq(W_{q}))-\mathcal{L}(W)\approx\frac{1}{2}(dq(W_{q})-W)^{\top}H(dq(W_{q})-W)$ (4)
+$\Delta\mathcal{L}$ $=\mathcal{L}(dq(W_{q}))-\mathcal{L}(W)$ (4)
+$\approx\frac{1}{2}\bigl(dq(W_{q})-W\bigr)^{\top}H\bigl(dq(W_{q})-W\bigr)$
 where $\mathcal{L}(W)$ is the loss function of the model evaluated at weights $W$.
@@
 (4) is minimized.
-Table 6 summarizes the computation strategies of $\delta$ adopted by various self-compensation quantization methods.
+Table VI summarizes the computation strategies of $\delta$ adopted by various self-compensation quantization methods.
+TABLE VI: Comparison of self-compensation quantization methods.
 Methods Weight perturbation $\delta$ computation
@@
 GQuant [15] Add gradient information
-Table 6: Comparison of self-compensation quantization methods.
 Some methods simplify the quantization problem to minimize the squared error between the quantized and original output of each layer [9, 10].
@@
 GQuant [15] further observes that the layer-wise MSE treats all hidden features equally, and therefore incorporates per-feature gradients into the objective function to improve quantization accuracy.
-Low-rank decomposition compensation
-Beyond compensating errors directly on the original weights, an alternative approach is to introduce an auxiliary branch dedicated to error correction.
+Low-rank compensation
+Beyond compensating errors directly on the original weights, an alternative approach is to introduce an auxiliary branch dedicated to quantization error mitigation.
 Some prior works leverage a low-rank term to reconstruct quantization error (e.g., ZeroQuant-v2 [39], CALDERA [29], LQER [41], QERA [42], etc), where the low-rank branch can either be quantized to higher precision, such as INT8, or remain in FP16.
@@
 The product $AB\in\mathbb{R}^{d\times k}$ serves as a low-rank approximation of the quantization error.
-Table 7 illustrates the computation process of the low-rank terms in some quantization methods.
+Table VII illustrates the computation process of the low-rank terms in some quantization methods.
+TABLE VII: Comparison of low-rank compensation quantization methods.A,BA,B represent the low-rank terms. $R_{XX}$ denotes the autocorrelation matrix of input $X$.
 Methods Low rank terms $A,B$ computation
-ZeroQuant-v2 [14] $\text{SVD}(dq(W_{q})-W)$
-LQER [16] $\text{SVD}(dq(W_{q})-W)$
-L2QER [16] $\text{SVD}(S(dq(W_{q})-W)),A:=S^{-1}A$
+ZeroQuant-v2 [39] $\mathrm{SVD}(dq(W_{q})-W)$
+LQER [41] $\mathrm{SVD}(dq(W_{q})-W)$
+L2QER [41] $\mathrm{SVD}(S(dq(W_{q})-W)),A:=S^{-1}A$
 CALDERA [29] LPLR factorization [29]
-QERA [22] $\text{SVD}(R^{\frac{1}{2}}_{X\!X}(dq(W_{q})-W)),A:=(R^{\frac{1}{2}}_{X\!X})^{-1}A$
-Table 7: Comparison of low-rank decomposition compensation quantization methods. $A,B$ represents the low rank terms. $R_{X\!X}$ denotes the autocorrelation matrix of the input $X$.
-ZeroQuant-v2 [14] and LQER [16] apply SVD composition to reconstruct the quantization error matrix $E=W-W_{q}$.
+QERA [42] $\mathrm{SVD}(R_{XX}^{1/2}(dq(W_{q})-W)),$ $A:=(R_{XX}^{1/2})^{-1}A$
+ZeroQuant-v2 [39] and LQER [41] apply SVD to reconstruct the quantization error matrix $E=W-dq(W_{q})$.
 Based on LQER, L2QER, which is proposed in LQER, scales the $E$ before applying SVD heuristically, and inversely applies scales in $A_{k}$.
@@
 In contrast to approaches that introduce low-rank branches externally, QwT [11] integrates lightweight linear layers within each decoder layer block of LLMs, directly compensating quantization errors, offering new insights for quantization error mitigation.
-2.2 Settings of quantization
-2.2.1 Symmetric vs. asymmetric
-Uniform quantizers can be broadly categorized into symmetric and asymmetric schemes, depending on how the zero point is handled during the mapping from floating-point to integer values.
-(a) Symmetric quantization
-(b) Asymmetric quantization
+II-B Settings of quantization
+II-B1 Symmetric vs. asymmetric
+Uniform quantizers can be broadly categorized into symmetric and asymmetric schemes, depending on how the zero point is handled during the mapping from floating-point to integer values, as illustrated in Fig.
+3.
+((a))
+((b))
 Figure 3: Symmetry of quantization
@@
 Symmetric quantization enables a very simple dequantization process.
-For linear layers, dequantization can be efficiently performed by multiplying the quantized output by the product of the scale factors for the weights and activations This simplifies computation and is often preferred in hardware implementations.
+For linear layers, dequantization can be efficiently performed by multiplying the quantized output by the product of the scale factors for the weights and activations.
+This simplifies computation and is often preferred in hardware implementations.
 However, symmetric quantization may lead to suboptimal representation when the data distribution is not centered around zero.
@@
 As such, the selection of quantization symmetry should be made with consideration of the specific use case and performance-complexity trade-offs.
-2.2.2 Granularity of quantization
+II-B2 Granularity of quantization
 The granularity of quantization often largely affects the low-bit quantization performance.
@@
 Inspired by this, subsequent papers like FlatQuant [34], ResQ [31], and OSTQuant [13] have also adopted per-channel granularity for quantization.
-2.3 FP4 Quantization and Variant Formats
+II-C FP4 Quantization and Variant Formats
 Recently, NVIDIA’s GeForce RTX 50 Series based on NVIDIA’s Blackwell architecture [25] has been released, which has enabled very low-precision operations using FP4 variants like MXFP4 [28, 26] and NVFP4 [24].
@@
 Fig.
-5 shows the weight distribution under BF16, INT4, and MXFP4.
+5 shows the weight distribution under BF16, INT4, MXFP4 and NVFP4.
 A clear observation is that under INT4 quantization, values with larger absolute magnitudes represent fewer data points, whereas MXFP4 more fully utilizes the representational capacity of 4-bit precision.
 Therefore, FP4 is more suitable than INT4 as a quantization data type.
-(a) Full-precision weight
-(b) INT4 quant weight
-(c) MXFP4 quant weight
+((a))
+((b))
+((c))
+((d))
 Figure 5: Weight distribution of a q_proj layer of LlaMA-3.2-1B.
@@
 (c): Weight distribution after MXFP4 quantization.
+(d): Weight distribution after NVFP4 quantization.
 LLM-FP4 [21] studies FP4 precision quantization for LLMs, including searching for the optimal exponent bias and maximal quantization value.
@@
 FP4-Quantization [36] improves the FP4 data representation while introducing a method called LREC (low-rank error calibration) to quantize weights into FP4.
-3 Extensive Evaluations
-Quantization settings We conduct experiments using W4A4 precision, comparing the results with the original precision (BF16 or FP16).
+III Extensive Evaluations
+Evaluation Scope.
+Based on the previous review and the defined experiment scope, we conducted extensive evaluations to systematically investigate the effects of Post-Training Quantization (PTQ) on large language models.
+Our analysis is organized into three main parts.
+In Section III-A, we evaluate several popular methods within the two-step quantization framework, providing a baseline for comparison.
+Section III-B performs an ablation study across different quantization settings, examining how these choices influence model performance.
+Finally, Section III-C applies the evaluated methods to FP4 and its variants, further exploring the potential and limitations of FP4 quantization.
+Quantization settings.
+We conduct experiments using W4A4 precision, comparing the results with the original precision (BF16 or FP16).
 In the main results, we use per-token or per-group (group size=128) asymmetric quantization for activations.
@@
 Models and tasks.
-We conduct experiments on LLaMA-3.2-1B, LLaMA-3.2-3B, and Qwen-2.5-3B.
+We conduct experiments on LLaMA-3.2 (1B, 3B), LLaMA-3.1 (8B, 70B), and Qwen2.5-3B.
+Additionally, we perform experiments on the MoE model Qwen1.5-MoE-A2.7B to further extend our evaluation.
 We evaluate their perplexity on the WikiText-2 dataset [23] and assess their accuracy on several zero-shot evaluation tasks, including commonsense reasoning, mathematics, sentiment analysis, and question answering.
-Specifically, we utilize the following datasets for evaluation: MMLU [31], SST2 [32], MathQA [33], ARC‑C [34], HellaSwag [35], WinoGrande [36], BoolQ [37], LAMBADA [38], and PIQA [39].
+Specifically, we utilize the following datasets for evaluation: MMLU [12], SST2 [33], MathQA [2], ARC‑C [5], HellaSwag [40], WinoGrande [30], BoolQ [7], LAMBADA [27] and PIQA [4].
 Implementation.
@@
 The training of the learnable orthogonal and scaling matrices follows the OSTQuant [13] procedure, using 1,000 samples from WikiText-2 and optimized with RiemannAdam.
-For the low-rank decomposition, we set the rank k=32 and store the low-rank matrices in FP16.
+For the low-rank compensation, we set the rank k=32 and store the low-rank matrices in FP16.
 For the non-trainable scaling matrix, we follow the SmoothQuant [38] calibration scheme with $\alpha$=0.5.
-3.1 Overall performance
+III-A Methods evaluation under the two-step quantization framework
+We select two of the most widely used pre-quantization transformation techniques, scaling and rotation, and combine them with two quantization error mitigation methods: GPTQ and low-rank compensation.
+A series of experiments is conducted to evaluate these combinations.
+Table VIII presents the results of using optimized scaling and rotation, illustrating the performance of different method combinations.
+To assess the statistical robustness of these results, we repeated the experiments of Table VIII across five different random seeds.
+Fig.
+7 visualizes the resulting error bars, where each bar represents the standard deviation of the Wiki2-PPL across the runs.
+The relatively small error bars indicate that the performance of each method is consistent and stable, confirming that the observed differences in Table VIII are statistically reliable.
+Table 7 presents the results obtained with randomly initialized rotations and calibrated scaling factors.
+In Table IX, we further conduct experiments on different models.
+The experiments of LLaMA-3.2-1B, LLaMA-3.2-3B, LLaMA-3.1-8B, LLaMA-3.1-70B and Qwen2.5-3B use the optimized rotation and scaling, and combine them with GPTQ and an additional low-rank compensation.
+For the MoE model Qwen1.5-MoE-A2.7B, given that the key advantage of MoE models lies in the fact that the actual computational cost scales only with the number of activated experts and performs full optimization of rotation and scaling parameters and GPTQ would be computationally expensive in quantization process compared to models with comparable computational budgets, we opt for more resource-efficient strategies which only use random rotation and low-rank compensation during PTQ.
+TABLE VIII: Evaluation of Quantization Method Combinations, in which rotation or scaling is optimized, in terms of Perplexity on WikiText2 and averaged accuracy on nine Zero-Shot tasks across per-channel and per-group settings of LlaMA-3.2-1B.
+PT and QEM represent pre-quantization transformation and quantization error mitigation, respectively.
 Evaluate Methods Channel Group-128
-PT QEM Wiki2‑PPL($\downarrow$) 0‑shot9 Avg.($\uparrow$) Wiki2‑PPL($\downarrow$) 0‑shot9 Avg.($\uparrow$)
+PT QEM Wiki2 PPL($\downarrow$) 0‑shot9 Avg.($\uparrow$) Wiki2 PPL($\downarrow$) 0‑shot9 Avg.($\uparrow$)
 BF16 9.76 54.78 9.76 54.78
@@
 Rotation+Scaling GPTQ+Low-rank 11.73 49.06 10.87 51.83
-Table 8: Evaluation of Quantization Method Combinations, in which rotation or scaling is optimized, in terms of Perplexity on WikiText2 and averaged accuracy on nine Zero-Shot tasks across per-channel and per-group settings of LlaMA-3.2-1B.
+Figure 6: Evaluation of Quantization Method Combinations, in which rotation or scaling is non-optimized, in terms of Perplexity on WikiText2 and averaged accuracy on nine Zero-Shot tasks across per-channel and per-group settings of LlaMA-3.2-1B.
 PT and QEM represent pre-quantization transformation and quantization error mitigation, respectively.
+Figure 7: Error bars showing the statistical robustness of the methods in Table VIII, computed over five runs with different random seeds.
 Evaluate Methods Channel Group-128
-PT QEM Wiki2‑PPL($\downarrow$) 0‑shot9 Avg.($\uparrow$) Wiki2‑PPL($\downarrow$) 0‑shot9 Avg.($\uparrow$)
+PT QEM Wiki2 PPL($\downarrow$) 0‑shot9 Avg.($\uparrow$) Wiki2 PPL($\downarrow$) 0‑shot9 Avg.($\uparrow$)
 BF16 9.76 54.78 9.76 54.78
@@
 Rotation+Scaling GPTQ+Low‑rank 13.91 46.46 12.17 49.41
-Table 9: Evaluation of Quantization Method Combinations, in which rotation or scaling is non-optimized, in terms of Perplexity on WikiText2 and averaged accuracy on nine Zero-Shot tasks across per-channel and per-group settings of LlaMA-3.2-1B.
-PT and QEM represent pre-quantization transformation and quantization error mitigation, respectively.
-We select two of the most widely used pre-quantization transformation techniques, scaling and rotation, and combine them with two quantization error mitigation methods: GPTQ and low-rank decomposition compensation.
-A series of experiments is conducted to evaluate these combinations.
-Table 8 presents the results of using optimized scaling and rotation, illustrating the performance of different method combinations.
-Table 9 presents the results obtained with randomly initialized rotations and calibrated scaling factors.
-Our experiments reveal that the best performance is achieved when both optimized scaling and rotation are applied in conjunction with GPTQ and the low-rank decomposition.
+TABLE IX: Evaluation of performance under different models.
+Metric Llama- 3.2-1B Llama- 3.2-3B Qwen2.5- 3B Qwen1.5- MoE-A2.7B Llama- 3.1-8B Llama- 3.1-70B
+FP16 W4A4 FP16 W4A4 FP16 W4A4 FP16 W4A4 FP16 W4A4 FP16 W4A4
+WikiText2 PPL($\downarrow$) 9.76 11.73 7.82 8.90 8.03 9.22 7.22 8.88 6.24 7.23 2.81 4.41
+Zero-Shot Accuracy($\uparrow$) General MMLU 36.76 29.21 54.06 46.70 65.13 56.32 60.86 51.38 63.32 56.98 75.26 70.57
+Sentiment SST2 66.86 55.28 73.62 68.58 90.37 69.15 67.89 60.32 77.18 59.52 82.11 75.69
+Math MathQA 28.94 25.43 34.47 30.55 37.15 32.50 35.01 29.21 39.50 34.91 51.66 45.36
+Reasoning PiQA 74.65 69.15 77.48 73.99 78.73 74.16 80.69 76.39 81.18 78.51 84.22 81.12
+HellaS 63.67 57.93 73.58 69.54 73.55 67.52 77.30 72.08 78.92 76.05 85.03 83.37
+WinoG 60.38 56.43 69.46 65.43 68.75 62.12 69.38 62.59 73.4 68.9 79.79 76.24
+QA BoolQ 63.67 59.02 72.97 68.75 77.58 65.41 79.60 73.98 82.05 78.75 85.5 83.39
+Lambada 62.10 56.78 70.08 65.32 66.80 57.77 71.38 61.69 75.33 72.48 78.75 76.75
+Arc-C 36.01 32.34 46.08 41.55 47.35 42.83 44.37 39.76 53.75 48.29 64.85 59.98
+Average 54.78 49.06 63.53 58.93 67.27 58.64 65.16 58.60 69.40 63.82 76.35 72.50
+In the following subsections, we provide a comprehensive and systematic analysis of the experimental results from multiple perspectives.
+III-A1 Pre-quantization transformation
+First, we analyze the performance of different pre-quantization transformation methods.
+Under optimization, using both optimized rotation and optimized scaling yields the best performance.
+As shown in Table VIII, applying optimized scaling together with optimized rotation always yields better performance.
+This is mainly because, after optimization, scaling can effectively fine-tune the inter-channel magnitude distribution of the rotated tensor, further improving the results and achieving better spatial utilization.
+Using only rotation also achieves significantly better performance compared with random rotation, which is consistent with the observation in SpinQuant [22] that optimization helps eliminate the high variance introduced by random rotations.
+Without optimization, it is better to use random rotation only.
+When the model is large, performing optimization is often impractical; therefore, we can only rely on random rotations and scaling factors obtained through calibration.
+From Table 7, we can see that in the absence of optimization, rotation alone performs better than rotation combined with scaling.
+We believe this is because when the scaling is applied after rotation, we hope it can fine-tune the rotated matrix.
+However, the calibration method becomes inaccurate in this case because outliers are averaged across channels, and the calculation of each channel’s scaling parameter cannot simply be based on its maximum value.
+III-A2 Quantization error mitigation
+Comparison between self-compensation and low-rank compensation.
+Between the two quantization error mitigation methods, we found that, except for the per-channel quantization case using only scaling, the effectiveness of low-rank compensation is often inferior to GPTQ.
+Under the optimal pre-quantization transformation setting (optimized rotation and scaling), using GPTQ alone results in perplexity of 11.8, whereas using only the low-rank method yields 14.58, under the per-channel settings.
+We believe the reason is that the low-rank method used in the experiments performs SVD based on the weight quantization error.
+It cannot accurately approximate the loss error, unlike methods like GPTQ that directly analyze the loss error.
+For low-rank compensation methods starting from the loss error, better results might be achieved, but it would require a relatively complex computational process, which we do not evaluate here.
+Combination of self-compensation and low-rank compensation.
+When both methods are used together, performance improves further, from 11.8 to 11.73 in the per-channel setting, and from 10.89 to 10.87 in the per-group setting, narrowing the gap to the full-precision model.
+We believe this is because low-rank compensation on the weight quantization error after GPTQ further compensates for the quantization error and leads to better results.
+III-A3 Relationship between the two-step
+The performance contributions of these two steps are additive and mutually complementary.
+Our experiments reveal that the best performance is achieved when both optimized scaling and rotation are applied in conjunction with GPTQ and the low-rank compensation.
 This configuration yields a perplexity of 11.73 in per-channel quantization for the Llama-3.2-1B model, which is close to the BF16 baseline of 9.76.
-Between the two quantization error mitigation methods, GPTQ clearly plays a more significant role than low-rank adaptation.
-Under the optimal pre-quantization transformation setting (optimized rotation and scaling), using GPTQ alone results in perplexity of 11.8, whereas using only the low-rank method yields 14.58, under the per-channel settings.
-When both methods are used together, performance improves further, from 11.8 to 11.73 in the per-channel setting, and from 10.89 to 10.87 in the per-group setting, narrowing the gap to the full-precision model.
-In addition, using both optimized rotation and scaling improves performance over using only optimized rotation.
-However, when employing non-optimized rotation and scaling, rotation alone tends to yield better results than the combination, which indicates that calibrated scaling may not perform well together with rotation in W4A4 settings.
-Model precision
-Wiki2
-PPL($\downarrow$)
-Zero-Shot Accuracy($\uparrow$)
-General Sentiment Math Reasoning QA Avg.
-MMLU SST2 MathQA PiQA HellaS WinoG BoolQ Lambada Arc-C
-Llama-
-3.2-1B
-fp16 9.76 36.76 66.86 28.94 74.65 63.67 60.38 63.67 62.1 36.01 54.78
-w4a4 11.73 29.21 55.28 25.43 69.15 57.93 56.43 59.02 56.78 32.34 49.06
-Llama-
-3.2-3B
-fp16 7.82 54.06 73.62 34.47 77.48 73.58 69.46 72.97 70.08 46.08 63.53
-w4a4 8.9 46.7 68.58 30.55 73.99 69.54 65.43 68.75 65.32 41.55 58.93
-Qwen-
-2.5-3B
-fp16 8.03 65.13 90.37 37.15 78.73 73.55 68.75 77.58 66.8 47.35 67.27
-w4a4 9.22 56.32 69.15 32.5 74.16 67.52 62.12 65.41 57.77 42.83 58.64
-Table 10: Comparison of symmetric and asymmetric quantization in Terms of Perplexity on WikiText2 and averaged accuracy on Zero-Shot tasks of LlaMA-3.2-1B
-In Table 10, we further conduct experiments on different models: LLaMA-3.2-1B, LLaMA-3.2-3B, and Qwen-2.5-3B, using the optimized scaling and rotation, and combined them with GPTQ and an additional low-rank decomposition.
-3.2 Symmetric vs. asymmetric
-Granularity Metric rotation+scaling+GPTQ Rotation+Scaling+GPTQ+Low-rank
-Sym W-Asym A-Asym Asym Sym W-Asym A-Asym Asym
-channel Wiki2-PPL($\downarrow$) 12.73 12.7 11.8 11.7 12.63 12.65 11.73 11.64
-Arc-Challenge Score(%) 31 31.7 33.3 33.7 31.9 29.9 32.3 32.7
-Table 11: Comparison of symmetric and asymmetric quantization in terms of Perplexity on WikiText2 and Arc [34] accuracy of LlaMA-3.2-1B
-In Table 11, we evaluate the performance of applying symmetric or asymmetric quantization to weights and activations, respectively.
+III-A4 Inference cost analysis
+Table X presents the theoretical analysis of inference cost for the evaluated methods.
+For experiments in Table VIII and 7, we only apply online rotation to the down-projection’s input in the MLP in each block.
+According to SpinQuant [22], which also applies online Hadamard rotation after ROPE to enhance key-value cache quantization performance, in addition to the inputs of the down-projection.
+And SpinQuant [22] has tested that these two online Hadamard processes introduce only a modest $\sim$8% increase in latency.
+Besides, the overhead of low-rank compensation is the most noteworthy aspect to discuss.
+As shown in Table X, we compute the additional memory cost and computation (FLOPs) of the low-rank computation.
+In our experiments, we set $r=32$.
+For Llama-3.2-1B, the low-rank compensation introduces approximately 1.41 M additional parameters per decoder layer, compared to about 61 M parameters in all linear layers, costing roughly 2.3% memory overhead additionally in FP16.
+After quantization, it costs roughly 9.2% memory overhead additionally.
+And it has an overhead computation ratio of $3.1\%$ in FP16.
+TABLE X: Inference Cost Analysis.
+Method Implementation Inference Overhead
+Scaling Fused into weight None
+Rotation Most: Fused into weight Fused rotation: None
+Few: Online Hadamard rotation Online rotation: Fast Hadamard
+GPTQ Directly weight changed None
+Low-rank Additional memory: $\begin{aligned} &A\in\mathbb{R}^{c_{\text{in}}\times r}\\ &B\in\mathbb{R}^{r\times c_{\text{out}}}\end{aligned}$ Memory: $\begin{aligned} \text{ratio}=\frac{r\times(c_{\text{in}}+c_{\text{out}})}{c_{\text{in}}c_{\text{out}}}\end{aligned}$
+Additional computation: XABXAB Computation: $\begin{aligned} \text{ratio}=\frac{r\times(c_{\text{in}}+c_{\text{out}})}{c_{\text{in}}\times c_{\text{out}}}\end{aligned}$
+III-B Evaluation of quantization settings
+III-B1 Symmetric vs. asymmetric
+TABLE XI: Comparison of symmetric and asymmetric quantization in terms of Perplexity on WikiText2 and Arc-challenge [5] accuracy of LlaMA-3.2-1B
+Granularity Metric Rotation+Scaling+GPTQ
+Sym W-Asym A-Asym Asym
+Channel Wiki2-PPL($\downarrow$) 12.73 12.7 11.8 11.7
+Arc-C Score(%) 31 31.7 33.3 33.7
+Granularity Metric Rotation+Scaling+GPTQ+Low-rank
+Sym W-Asym A-Asym Asym
+Channel Wiki2-PPL($\downarrow$) 12.63 12.65 11.73 11.64
+Arc-C Score(%) 31.9 29.9 32.3 32.7
+In Table XI, we evaluate the performance of applying symmetric or asymmetric quantization to weights and activations, respectively.
 The results demonstrate that asymmetric quantization for activations leads to a significant improvement over symmetric quantization, while the performance gain for weights is relatively minor.
@@
 This setup not only achieves much better performance than fully symmetric quantization, but also reduces the computational overhead associated with the dequantization process, compared to fully asymmetric quantization for both weights and activations.
-3.3 Granularity
+III-B2 Granularity
 As illustrated in Fig.
-6, we conduct a comprehensive comparison of per-group quantization performance across varying group sizes, measured by perplexity on WikiText2.
+XII, we conduct a comprehensive comparison of per-group quantization performance across varying group sizes, measured by perplexity on WikiText2.
 The experimental results reveal that finer quantization granularity yields superior model performance (lower perplexity), albeit at the expense of elevated storage requirements for quantization parameters.
@@
 Quantitative analysis of these additional bit costs is presented in Table 8, which compares the average bit overhead introduced by scaling factors under symmetric weight quantization.
-Figure 6: Comparison of different group sizes of quantization in terms of Perplexity on WikiText2 and averaged accuracy on Zero-Shot tasks of LlaMA-3.2-1B
+Figure 8: Average extra bits introduced by FP32 scaling factors of different granularities
+TABLE XII: Comparison of different group sizes of quantization in terms of Perplexity on WikiText2 and averaged accuracy on Zero-Shot tasks of LlaMA-3.2-1B
 Granularity group-32 group-64 group-128 group-256 group-512
-Average Extra bits
-1 0.5 0.25 0.125 0.0625
-Table 12: Average extra bits introduced by FP32 scaling factors of different granularities
-3.4 FP4 vs. INT4
+Average Extra bits 1 0.5 0.25 0.125 0.0625
+III-C Extending Experiments in FP4
+III-C1 FP4 vs. INT4
 Extending our experiments to the FP4 format, we compare several variants of FP4, including MXFP4, NVFP4, and naive FP4 implementation with varying granularities and scaling factor formats.
@@
 To ensure a fair comparison with MXFP4 and NVFP4, all other experiments using per-group quantization adopt symmetric quantization for both activations and weights.
+TABLE XIII: Evaluation of INT4 and FP4 format quantization without using pre-quantization transformation, in terms of Perplexity on WikiText2 of LlaMA-3.2-1B.
+Channel means quantization granularity is per-channel; g16 and g32 denote per-group quantization with group sizes of 16 and 32, respectively.
+FP32, E8M0, and E4M3 refer to the format of the scaling factor.
+And for E4M3, the scaling strategy follows NVFP4’s approach, employing a two-level scaling scheme.
+Data Format INT4 channel, FP32 FP4 channel, FP32 INT4 g16, FP32 FP4 g16, FP32
+RTN 256.97 135.68 11.75 11.34
+GPTQ 161.04 101.36 11.47 10.87
+Low-rank 179.87 132.14 11.69 11.32
+GPTQ+Low-rank 139.31 96.33 10.92 10.76
+Data Format MXFP4 g32, E8M0 FP4 g32, E4M3 FP4 g16, E8M0 NVFP4 g16, E4M3
+RTN 15.17 12.04 14.52 11.54
+GPTQ 13.24 11.36 12.77 10.98
+Low-rank 15.56 11.98 14.67 11.48
+GPTQ+Low-rank 12.69 11.18 12.25 10.83
 First, we apply no extra pre-quantization transformation, conducting experiments using only quantization error mitigation methods.
-Table 13 demonstrates that FP4 (per-channel) shows much better performance than INT4 (per-channel).
+FP4 performs better than INT4 without pre-quantization transformation in per-channel setting.
+Table XIII demonstrates that FP4 (per-channel) shows much better performance than INT4 (per-channel).
 We believe this is because, in the absence of the pre-quantization transformation, FP4 better aligns with the data distribution, thus achieving superior results.
-For finer granularity scenarios, FP4 and INT4 show very small differences, and both the format of the scaling factor and the size of the group significantly impact performance.
+For finer granularity scenarios, FP4 and INT4 show very small differences.
+Table XIII demonstrates that FP4 and INT4 show minimal differences at finer granularities, with FP4 performing only slightly better than INT4.
+We attribute this to the fact that fine-grained quantization helps mitigate the impact of outliers.
+Consequently, the advantage of FP4 in handling outliers is reduced under finer granularity, resulting in only a marginal performance improvement over INT4.
 Compared to MXFP4, NVFP4 achieves better performance due to its smaller group size and more precise scaling factor format.
+We believe that both the format of the scaling factor and the size of the group significantly impact performance.
 With the E4M3 variant of FP8, NVFP4 can encode scaling factors with fractional precision.
@@
 NVFP4 performs better than MXFP4, showing a lower risk of noticeable accuracy drop, matching NVIDIA’s official conclusions [1].
-Quantization
-error
-mitigation
-Data Format
-INT4
-channel, FP32
-FP4
-channel, FP32
-INT4
-g16, FP32
-FP4
-g16, FP32
-MXFP4
-g32, E8M0
-FP4
-g32, E4M3
-FP4
-g16, E8M0
-NVFP4
-g16, E4M3
-RTN 256.97 135.68 11.75 11.34 15.17 12.04 14.52 11.54
-GPTQ 161.04 101.36 11.47 10.87 13.24 11.36 12.77 10.98
-low-rank 179.87 132.14 11.69 11.32 15.56 11.98 14.67 11.48
-GPTQ+low-rank 139.31 96.33 10.92 10.76 12.69 11.18 12.25 10.83
-Table 13: Evaluation of INT4 and FP4 format quantization without using pre-quantization transformation, in terms of Perplexity on WikiText2 of LlaMA-3.2-1B.
+TABLE XIV: Evaluation of INT4, MXFP4, and NVFP4 using both pre-quantization transformation and quantization error mitigation techniques, in terms of Perplexity on WikiText2 of LlaMA-3.2-1B.
 Channel means quantization granularity is per-channel; g16 and g32 denote per-group quantization with group sizes of 16 and 32, respectively.
 FP32, E8M0, and E4M3 refer to the format of the scaling factor.
-And for E4M3, the scaling strategy follows NVFP4’s approach, employing a two-level scaling scheme.
-Quantization
-error
-mitigation
-Data Format
-NVFP4
-W:E4M3, A:E4M3
-FP4
-W:E2M1, A:E4M3
-FP4
-W:E2M1, A:E2M1
-FP4
-W:INT4, A:E4M3
-FP4
-W:INT4, A:INT4
+Evaluate Methods Wiki2‑PPL($\downarrow$)
+Pre-quantization transformation Quantization error mitigation INT4 channel, FP32 FP4 channel, FP32 MXFP4 g32, E8M0 NVFP4 g16, E4M3
+Scaling GPTQ 134.85 89.22 13.24 10.97
+Scaling Low-rank 165.4 104.55 14.91 11.36
+Scaling GPTQ+Low-rank 124.56 114.96 12.58 10.81
+Optimized rotation+scaling GPTQ 11.8 12.13 12.37 11.04
+Optimized rotation+scaling Low-rank 14.58 15.15 15.43 12.57
+Optimized rotation+scaling GPTQ+Low-rank 11.73 12.13 12.29 11.01
+Then, we apply pre-quantization transformation together with quantization error mitigation on FP4 and main FP4 variants to compare with INT4, and the results are shown in Table XIV.
+Rotation-based methods yield smaller improvements for MXFP4 and NVFP4 than for INT4 (per-channel).
+As argued in AMXFP4 [16], grouping activation tensors into small micro-scaled units mitigates outliers like rotation.
+For MXFP4 and NVFP4 formats, the group size is extremely small, so data rotation contributes little and may even lead to degradation.
+FP4 does not show its advantage under rotation.
+Although MXFP4 and NVFP4 perform better than INT4 (per-channel), it is observed that under the same channel-wise setting, INT4 performs better than FP4 after rotation, while when applying only scaling, FP4 still outperforms INT4.
+We believe this is because rotation can largely eliminate outliers, to the extent that the advantage of FP4 over INT4 when dealing with outliers diminishes.
+Consequently, INT4 demonstrates stronger performance in handling more uniform distributions.
+III-C2 Exploration of FP4 scaling factor
+In the analysis of Table XIII in Section III-C1, we observed that the scaling factor has a significant impact, and we conduct further exploratory research on this aspect.
+We further reduced the bit-width of the scaling factor to 4 bits to evaluate its effectiveness under such a low-bit setting.
+As shown in Table XV, applying 4-bit scaling factors to both weights and activations leads to very poor performance.
+When applied only to weights, the performance remains moderate but is worse compared to NVFP4.
+TABLE XV: Evaluation of FP4 format quantization using low bit scaling factor and group size 16, in terms of Perplexity on WikiText2 of LlaMA-3.2-1B.
+E4M3, E2M1, and INT4 refer to the format of the scaling factor.
+W and A denote weight and activation, respectively.
+All of these follow NVFP4’s approach, employing a two-level scaling scheme.
+Quantization error mitigation Data Format
+NVFP4 W:E4M3 A:E4M3 FP4 W:E2M1 A:E4M3 FP4 W:E2M1 A:E2M1 FP4 W:INT4 A:E4M3 FP4 W:INT4 A:INT4
 RTN 11.54 14.10 4130 14.66 4777
@@
 GPTQ+Low-rank 10.83 11.38 3650 11.59 4124
-Table 14: Evaluation of FP4 format quantization using low bit scaling factor and group size 16, in terms of Perplexity on WikiText2 of LlaMA-3.2-1B.
-E4M3, E2M1, and INT4 refer to the format of the scaling factor.
-W and A denote weight and activation, respectively.
-All of these follow NVFP4’s approach, employing a two-level scaling scheme.
-We further reduce the bit-width of the scaling factor to 4 bits to evaluate its effectiveness under such a low-bit setting.
-As shown in Table 14, applying 4-bit scaling factors to both weights and activations leads to very poor performance.
-When applied only to weights, the performance remains moderate and is worse compared with NVFP4.
-Evaluate Methods Wiki2‑PPL($\downarrow$)
-Pre-quantization
-transformation
-Quantization error
-mitigation
-INT4
-channel, FP32
-MXFP4
-g32, E8M0
-NVFP4
-g16, E4M3
-scaling GPTQ 134.85 13.24 10.97
-scaling low-rank 165.4 14.91 11.36
-scaling GPTQ+low-rank 124.56 12.58 10.81
-optimized rotation+scaling GPTQ 11.8 12.37 11.04
-optimized rotation+scaling low-rank 14.58 15.43 12.57
-optimized rotation+scaling GPTQ+low-rank 11.73 12.29 11.01
-Table 15: Evaluation of INT4, MXFP4, and NVFP4 using both pre-quantization transformation and quantization error mitigation techniques, in terms of Perplexity on WikiText2 of LlaMA-3.2-1B.
-Channel means quantization granularity is per-channel; g16 and g32 denote per-group quantization with group sizes of 16 and 32, respectively.
-FP32, E8M0, and E4M3 refer to the format of the scaling factor.
-Notably, rotation-based methods yield smaller improvements for MXFP4 and NVFP4 than for INT4 (per-channel).
-As argued in AMXFP4 [41], grouping activation tensors into small micro-scaled units mitigates outliers like rotation.
-For MXFP4 and NVFP4 formats, the group size is extremely small, so data rotation contributes little and may even lead to degradation.
-4 Conclusions
+IV Conclusions
 In this paper, we have presented an experimental evaluation of Large language model quantization.
 We systematically decouple the commonly used quantization into two key steps: pre-quantization transformation and quantization error mitigation.
-After a comprehensive review of existing approaches, we choose scaling and rotation in pre-quantization transformation step, GPTQ, and low-rank compensation in the quantization error mitigation step to further conduct experimental evaluations on the same ground.
+After a comprehensive review of existing approaches, we choose scaling and rotation in the pre-quantization transformation step, GPTQ, and low-rank compensation in the quantization error mitigation step to further conduct experimental evaluations on the same ground.
 We found that using the optimized rotation and scaling, together with two compensation methods, demonstrates the best performance.
 For quantization settings, asymmetric quantization and smaller group sizes yield better results.
-And using asymmetric quantization on activations shows more benefits on performance compared with weights.
+And using asymmetric quantization on activations shows more benefits in performance compared with weights.
 For FP4 quantization, we can adopt methods used in INT4 quantization.
-However, rotation-based pre-quantization transformation yields no observable more performance gains for MXFP4 and NVFP4 than scaling-based methods.
+However, rotation-based pre-quantization transformation yields no observable performance gains for MXFP4 and NVFP4 than scaling-based methods.
 And we find that the format of scaling factors demonstrates a large effect.